### v1

# The intrinsic subgroup of an elliptic curve and Mazur’s torsion theorem

Abstract.
We define and study a biadditive symmetric (not necessarily perfect) pairing on the torsion part $\mathrm{Pic}(X)_{{\operatorname{tors}}}$ of the Picard group of a smooth projective curve $X$ over a field $k$ with values in $k^{\times}\otimes\mathbb{Q}/\mathbb{Z}$. We call its kernel the intrinsic subgroup of $X$. It turns out that some information on the reduction type of $X$ can be read off from the intrinsic subgroup. Mazur’s torsion theorem says that there are exactly 15 isomorphism classes of abelian groups that appear as the rational torsion points of an elliptic curve $X$ over $\mathbb{Q}$ (identified with $\mathrm{Pic}(X)_{\operatorname{tors}}$). We refine this result by determining which subgroups of those 15 groups appear as the intrinsic subgroups.
Key words and phrases:
Picard group of a curve, torsion rational points of an elliptic curve.
2010 Mathematics Subject Classification:
11G18, 14G05, 14G35
2010 Mathematics Subject Classification:
11G05 (11G18, 14G25, 14G35)
1. Introduction
Let $X$ be a geometrically irreducible smooth projective curve over a field $k$. In §2, we shall construct a biadditive symmetric pairing
(1.1) $\langle\cdot,\cdot\rangle:\operatorname{Pic}(X)_{\operatorname{tors}}\times\operatorname{Pic}(X)_{\operatorname{tors}}\to k^{\times}\otimes\mathbb{Q}/\mathbb{Z},$
where $\operatorname{Pic}(X)_{\operatorname{tors}}$ is the torsion part of the Picard group $\operatorname{Pic}(X)$ of $X$. We then define the intrinsic subgroup of $\operatorname{Pic}(X)_{\operatorname{tors}}$ by
$\operatorname{Pic}(X)_{\operatorname{tors}}^{\operatorname{is}}:=\{a\in\operatorname{Pic}(X)_{\operatorname{tors}}\mid\langle a,b\rangle=0\text{ for all }b\in\operatorname{Pic}(X)_{\operatorname{tors}}\}.$
We remark that (1.1) is not entirely new. Indeed, its finite coefficient analogue (with values in $k^{\times}/(k^{\times})^{m}$) is considered by Frey and Rück in [4] (see Remark 2.2). The first and second authors encountered with a disguised version of (1.1) in [18] (see Remark 2.3). Obviously, (1.1) is trivial if $k^{\times}\otimes\mathbb{Q}/\mathbb{Z}=0$, e.g. when $k$ is finite, algebraically closed, or $k=\mathbb{R}$. On the other hand, $\operatorname{Pic}(X)_{\operatorname{tors}}^{\operatorname{is}}$ appears to be a non-trivial new invariant if $k$ is a global field or a $p$-adic field.
We will mainly consider the case where $X=E$ is an elliptic curve so that we may identify $E(k)_{\operatorname{tors}}=\operatorname{Pic}(E)_{\operatorname{tors}}$ by $P\mapsto[P-O]$, where $O\in E(k)$ is the identity element. Thus we write $\langle P,Q\rangle:=\langle[P-O],[Q-O]\rangle$ for $P,Q\in E(k)_{\operatorname{tors}}$ and $E(k)_{\operatorname{tors}}^{\operatorname{is}}:=\operatorname{Pic}(E)_{\operatorname{tors}}^{\operatorname{is}}$. Our main result is a classification of possible structures of $(E(\mathbb{Q})_{\operatorname{tors}},E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}})$ for an elliptic curve $E$ over $\mathbb{Q}$. A celebrated theorem of Mazur [10, Theorem (8)] states that $E(\mathbb{Q})_{\operatorname{tors}}$ is either cyclic of order $1,\ldots,10$, or $12$, or is isomorphic to $(\mathbb{Z}/2N\mathbb{Z})\times(\mathbb{Z}/2\mathbb{Z})$, $N=1,2,3,4$, and all cases are realized by infinitely many (mutually non-isomorphic) elliptic curves. The classification of $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}$ is given in the following theorem, whose proof will occupy §36.
Theorem 1.1.
Let $E$ be an elliptic curve over $\mathbb{Q}$. Then there is a pair $(A,B)$ of an abelian group $A$ and its subgroup $B$ in the following table, together with an isomorphism $\alpha:E(\mathbb{Q})_{\operatorname{tors}}\cong A$, such that $\alpha(E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}})=B$.
[…]
(For $A=\mathbb{Z}/4\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$, there are two $\operatorname{Aut}(A)$-conjugacy classes of order two subgroups, both of which can be taken as $B$.) Moreover, for each possible pair $(A,B)$ as above, there are infinitely many (mutually non-isomorphic) elliptic curves $E$ over $\mathbb{Q}$ for which there is an isomorphism $\alpha:E(\mathbb{Q})_{\operatorname{tors}}\cong A$ such that $\alpha(E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}})=B$.
We briefly explain the outline of the proof. Our approach is completely explicit. As usual, we let $X_{1}(N)$ denote the modular curve associated to the congruence subgroup $\Gamma_{1}(N)$. It possesses a model over $\mathbb{Q}$ on which the cusp 0 is $\mathbb{Q}$-rational. For a subfield $k$ of $\mathbb{C}$, the non-cuspidal points in $X_{1}(N)(k)$ parameterize isomorphism classes of pairs $(E,P)$ of an elliptic curve $E$ over $k$ and a $k$-rational point $P$ of order $N$. More precisely, if the coordinates of a point $\tau\in\mathbb{H}$ on this model of $X_{1}(N)$ is $k$-rational, then the isomorphism class of $(\mathbb{C}/(\mathbb{Z}\tau+\mathbb{Z}),1/N)$ contains a pair $(E,P)$ of an elliptic curve $E$ over $k$ and a $k$-rational NN-torsion point $P$. A famous theorem of Mazur [10, Theorem (7)] says that $X_{1}(N)(\mathbb{Q})$ has a non-cuspidal point precisely when $N=1,\ldots,10$, or $N=12$.
For a pair $(N,M)$ of $N\in\{4,\dots,10,12\}$ and a positive divisor $M$ of $N$, we will construct a smooth projective curve $X_{1}(N)^{\pm}_{M}$ over $\mathbb{Q}$ equipped with a finite morphism $X_{1}(N)_{M}^{\pm}\to X_{1}(N)$ that enjoys the following property: when $t\in X_{1}(N)(\mathbb{Q})$ corresponds to a pair $(E,P)$ as above (with $k=\mathbb{Q}$), one has
$\langle P,(N/M)P\rangle=0\Leftrightarrow t\in\operatorname{Im}(X_{1}(N)^{\pm}_{M}(\mathbb{Q})\to X_{1}(N)(\mathbb{Q})).$
It is worth mentioning that $X_{1}(N)^{+}_{M}$ is isomorphic to the modular curve $X_{1}(MN,N)$ associated to $\Gamma_{1}(MN,N):=\Gamma_{0}(MN)\cap\Gamma_{1}(N)$ (Theorem 5.6) and $X_{1}(N)^{-}_{M}$ is its twist.
Similarly, for a positive even integer $N$, we let $X_{1}^{0}(N,2)$ denote the modular curve associated to the congruence subgroup
(1.2) $\Gamma_{1}^{0}(N,2):=\Gamma_{1}(N)\cap\Gamma^{0}(2)=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in{\operatorname{SL}}(2,\mathbb{Z}):a,d\equiv 1\bmod N,~~2|b,~N|c\right\}.$
The modular curve $X_{1}^{0}(N,2)$ parameterizes isomorphism classes of triples $(E,P,Q)$ of an elliptic curve $E$ and $k$-rational points $P,Q$ such that $P$ and $Q$ are of order $N$ and $2$ respectively, and $Q\notin\langle P\rangle$. For $N\in\{4,6,8\}$ and $\underline{M}=(M_{1},M_{2},M_{3})$ with $M_{1}|N,M_{2},M_{3}\in\{1,2\}$, we will construct a smooth projective curve $X_{1}^{0}(N,2)_{\underline{M}}^{\pm}$ over $\mathbb{Q}$ equipped with a finite morphism to $X_{1}^{0}(N,2)$, by which one can interpret the conditions
$\langle P,\frac{N}{M_{1}}P\rangle=\langle Q,\frac{2}{M_{2}}Q\rangle=\langle P,\frac{2}{M_{3}}Q\rangle=0$
for a triple $(E,P,Q)$ as above corresponding to a point of $X_{1}^{0}(N,2)(\mathbb{Q})$. Therefore Theorem 1.1 is reduced to a study of the $\mathbb{Q}$-rational points of $X_{1}(N)_{M}^{\pm}$ and $X_{1}^{0}(N,2)_{\underline{M}}^{\pm}$. As our construction of these curves are explicit, this can be done by a (more or less) direct computation.
In the last section §7, we generalize the pairing (1.1) and the intrinsic subgroup to higher dimensional varieties. This new construction enables us to prove that, if $X$ has good reduction with respect to a discrete valuation of $k$, there is a strong restriction on the values of the pairing (see Proposition 7.3). Finally, as a sample for the case of bad reduction, we compute the intrinsic subgroup of Tate elliptic curves (see Proposition 7.7). As an application, we show that the intrinsic subgroup of an elliptic curve over a number field imposes some constraint on its reduction type (see Corollary 7.8).
Acknowledgement
It is Kenneth Ribet who formulated (1.1) as a pairing and asked if it is symmetric. The authors express deep gratitude to him for asking this question, which was a starting point of our work.
The first author (T. Y.) is supported by JSPS KAKENHI Grant (25K06961). The second author (Y. Y.) is supported by Grant 113-2115-M-002-003-MY3 of the National Science and Technology Council of the Republic of China (Taiwan). The third author (H. Y.) is supported by National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2023-00239918 and No. 2020R1A5A1016126). The fourth author (M. Y.) is supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (RS-2025-23525445).
Notation and convention
For an abelian group $A$, we write $A[n]$ and $A/n$ for the kernel and cokernel of $A\to A,\ a\mapsto na$ for each $n\in\mathbb{Z}$, and put $A_{\operatorname{tors}}:=\cup_{n>0}A[n]$. For a field $k$, we set $\mu_{n}(k):=(k^{\times})[n]$ and $\mu(k):=(k^{\times})_{\operatorname{tors}}$.
The identity element of an elliptic curve is denoted by $O$.
2. Pairing on the Picard group and the intrinsic subgroup
Throughout this section, we let $X$ be a geometrically irreducible smooth projective curve over a field $k$. The goal of this section is to construct a biadditive pairing
(2.1) $\langle\cdot,\cdot\rangle:\operatorname{Pic}(X)_{\operatorname{tors}}\times\operatorname{Pic}^{0}(X)\to k^{\times}\otimes\mathbb{Q}/\mathbb{Z}$
which restricts to a symmetric pairing (1.1). Here $\operatorname{Pic}^{0}(X):=\ker(\deg:\operatorname{Pic}(X)\to\mathbb{Z})$ (thus $\operatorname{Pic}(X)_{\operatorname{tors}}\subset\operatorname{Pic}^{0}(X)$). This will be generalized to higher dimensional varieties in Section 7 below, but our construction in this section is more elementary and sufficient for most of our discussion.
2.1. Construction of the pairing
For a non-zero rational function $f\in k(X)^{\times}$, a closed point $Q\in X$, and $n\in\mathbb{Z}_{>0}$ such that $n|{\operatorname{ord}}_{Q}(f)$, we define the leading coefficient by
${\operatorname{lc}}_{Q}(f,n):=\left(\frac{f}{\pi^{{\operatorname{ord}}_{Q}(f)}}(Q)\right)\otimes\frac{1}{n}\in k(Q)^{\times}\otimes\mathbb{Q}/\mathbb{Z},$
where $\pi\in k(X)^{\times}$ is a uniformizer at $Q$. This is independent of the choice of $\pi$: replacing $\pi$ by $u\pi$ for some $u\in k(X)^{\times}$ with ${\operatorname{ord}}_{Q}(u)=0$ changes ${\operatorname{lc}}_{Q}(f,n)$ by $u(Q)^{-{\operatorname{ord}}_{Q}(f)}\otimes(1/n)$ which vanishes by the assumption $n|{\operatorname{ord}}_{Q}(f)$. Note also that for any $m>0$ we have
(2.2) ${\operatorname{lc}}_{Q}(f,n)={\operatorname{lc}}_{Q}(f^{m},mn).$
We define a subgroup $\operatorname{Div}_{t}(X)$ of the group of divisors $\operatorname{Div}(X)$ on $X$ by
$\operatorname{Div}_{t}(X):=\{D\in\operatorname{Div}(X)\mid[D]\in\operatorname{Pic}(X)_{\operatorname{tors}}\},$
where we write $D]\in\operatorname{Pic}(X)$ for the class of $D\in\operatorname{Div}(X)$. It contains all rationally trivial divisors, and all its members have degree zero:
$\{\operatorname{div}(f)\mid f\in k(X)^{\times}\}\subset\operatorname{Div}_{t}(X)\subset\operatorname{Div}^{0}(X):=\ker(\deg:\operatorname{Div}(X)\to\mathbb{Z}).$
Given $D\in\operatorname{Div}_{t}(X)$ and $E=\sum_{j}e_{j}Q_{j}\in\operatorname{Div}^{0}(X)$, we define
(2.3) $\langle D,E\rangle:=\prod_{j}N_{k(Q_{j})/k}({\operatorname{lc}}_{Q_{j}}(f,n))^{e_{j}}\in k^{\times}\otimes\mathbb{Q}/\mathbb{Z},$
where $n\in\mathbb{Z}_{>0}$ is such that $n[D]=0$, $f\in k(X)^{\times}$ is such that $\operatorname{div}(f)=nD$ (so that $n|{\operatorname{ord}}_{P}(f)$ for any $P$), and $N_{k(Q_{j})/k}:k(Q_{j})^{\times}\otimes\mathbb{Q}/\mathbb{Z}\to k^{\times}\otimes\mathbb{Q}/\mathbb{Z}$ is induced by the norm map attached to the finite extension $k(Q_{j})/k$.
Proposition 2.1.
For $D\in\operatorname{Div}_{t}(X)$ and $E\in\operatorname{Div}^{0}(X)$, the element (2.3) depends only on the classes of $D,E$ in $\operatorname{Pic}(X)$. Consequently we obtain the induced biadditive pairing (2.1) (denoted by the same symbol). Moreover, its restriction to $\operatorname{Pic}(X)_{\operatorname{tors}}\times\operatorname{Pic}(X)_{\operatorname{tors}}$ is symmetric.
Proof.
We first show that (2.3) is independent of the choices of $f$ and $n$.
  • Replacing $f$ by $cf$ for some $c\in k^{\times}$ does not change $\langle D,E\rangle$ since $\prod_{j}N_{k(Q_{j})/k}(c)^{e_{j}}=\prod_{j}c^{[k(Q_{j}):k]e_{j}}=c^{\deg(E)}=1$ by $\deg(E)=0$.
  • For any $m\in\mathbb{Z}_{>0}$, replacing $n$ by $mn$ does not change $\langle D,E\rangle$ since we have $\operatorname{div}(f^{m})=mnD$ and (2.2).
In particular, if $D=\operatorname{div}(f)$ for $f\in k(X)^{\times}$, we may take $n=1$ to conclude $\langle D,E\rangle=\langle\operatorname{div}(f),E\rangle=0$. Thus (2.3) factors through $\langle\cdot,\cdot\rangle:\operatorname{Pic}(X)_{\operatorname{tors}}\times\operatorname{Div}^{0}(X)\to k^{\times}\otimes\mathbb{Q}/\mathbb{Z}.$
We now assume that $E$ is also in $\operatorname{Div}_{t}(X)$ and that $|D|\cap|E|=\emptyset$. We claim
(2.4) $\langle D,E\rangle=\langle E,D\rangle.$
To show this, we suppose $n[D]=n[E]=0$ by replacing $n$ if necessary, and write $D=\sum_{i}d_{i}P_{i}$ and $nE=\operatorname{div}(g)$ with $g\in k(X)^{\times}$. Then we have ${\operatorname{lc}}_{Q}(f,n)=f(Q)\otimes(1/n)$ and ${\operatorname{lc}}_{P}(g,n)=g(P)\otimes(1/n)$ for all $P\in|D|$ and $Q\in|E|$, and hence
$\langle D,E\rangle=\prod_{j}N_{k(Q_{j})/k}(f(Q_{j}))^{e_{j}}\otimes\frac{1}{n},\qquad\langle E,D\rangle=\prod_{i}N_{k(P_{i})/k}(g(P_{i}))^{d_{i}}\otimes\frac{1}{n}.$
Recall that the Weil pairing ${\operatorname{Weil}}_{n}([D],[E])$ is given by
${\operatorname{Weil}}_{n}([D],[E])=\frac{\prod_{j}N_{k(Q_{j})/k}(f(Q_{j}))^{e_{j}}}{\prod_{i}N_{k(P_{i})/k}(g(P_{i}))^{d_{i}}}$
which takes its value in $\mu_{n}(k)$. (This is a consequence of the Weil reciprocity and valid in any characteristic, although ${\operatorname{Weil}}_{n}$ fails to be perfect if $p=\operatorname{char}(k)>0$ and $p|n$.) Hence it vanishes after taking $\otimes(1/n)$ in $k^{\times}\otimes\mathbb{Q}/\mathbb{Z}$. We have shown (2.4).
Finally, suppose that $E=\operatorname{div}(g)$ for some $g\in k(X)^{\times}$. We can always find $h\in k(X)^{\times}$ such that $|D+\operatorname{div}(h)|\cap|\operatorname{div}(g)|=\emptyset$. From what we have shown it follows that
$\langle D,\operatorname{div}(g)\rangle=\langle D+\operatorname{div}(h),\operatorname{div}(g)\rangle=\langle\operatorname{div}(g),D+\operatorname{div}(h)\rangle=0.$
This completes the proof of the proposition. ∎
Remark 2.2.
As is seen from the proof, there is an analogous biadditive pairing
$\langle\cdot,\cdot\rangle_{n}:\operatorname{Pic}(X)[n]\times\operatorname{Pic}^{0}(X)/n\to k^{\times}/(k^{\times})^{n}$
for each $n>0$, which satisfies $\langle d,e\rangle_{n}={\operatorname{Weil}}_{n}(d,e)\langle e,d\rangle_{n}$for any $d,e\in\operatorname{Pic}(X)[n]$. This pairing has been constructed by Frey and Rück in [4] when $k$ is a finite field and shown to be perfect if $|k|\equiv 1\bmod n$. (This is called the Lichtenbaum-Tate pairing in some literature, see e.g. [16, XI.9].)
Remark 2.3 (Relation with the generalized Jacobian).
Let $P_{0},\dots,P_{r}\in X(k)$ be distinct $k$-rational points on $X$ such that $P_{i}-P_{0}\in\operatorname{Div}_{t}(X)$ for any $i$, and put $D:=P_{0}+\cdots+P_{r}$. Let $J$ (resp. $\widetilde{J}$) be the Jacobian of $X$ (resp. generalized Jacobian of $X$ with modulus $D$). We have an exact sequence
$0\to T\to\widetilde{J}\to J\to 0,$
where $T\cong\mathbb{G}_{m}^{r}$ is the (split) torus with character group $\mathbb{Z}[D]^{0}:=\oplus_{i=1}^{r}(P_{i}-P_{0})\mathbb{Z}$. It induces a diagram with exact row:
(2.5) ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${T(k)_{\operatorname{tors}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\widetilde{J}(k)_{\operatorname{tors}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${J(k)_{\operatorname{tors}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi}$${H^{1}(k,T(\overline{k})_{\operatorname{tors}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathbb{Z}[D]^{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{[\cdot]}$${(k^{\times}\otimes\mathbb{Q}/\mathbb{Z})^{r},}$
where $\Phi$ is defined by the commutativity. It is shown in [18, Lemma 2.3.1] that $\Phi$ can be described in terms of the pairing (2.2) as
$\Phi([P_{i}-P_{0}])=(\langle P_{i}-P_{0},P_{j}-P_{0}\rangle)_{j=1}^{r}.$
2.2. The intrinsic subgroup
Definition 2.4.
We define the intrinsic subgroup of $\operatorname{Pic}(X)_{\operatorname{tors}}$ by
$\operatorname{Pic}(X)_{\operatorname{tors}}^{\operatorname{is}}:=\{a\in\operatorname{Pic}(X)_{\operatorname{tors}}\mid\langle a,b\rangle=0\text{ for all }b\in\operatorname{Pic}(X)_{\operatorname{tors}}\}.$
If $X=E$ is an elliptic curve, we identify $E(k)=\operatorname{Pic}^{0}(X)$ and put $E(k)_{\operatorname{tors}}^{\operatorname{is}}:=\operatorname{Pic}(X)_{\operatorname{tors}}^{\operatorname{is}}$.
Here we exhibit a few examples in the case of modular curves.
Example 2.5.
Let $p$ be a prime and consider the modular curve $X=X_{0}(p)$ over $k=\mathbb{Q}$. Let $P_{0}$ and $P_{1}$ be the 0 and $\infty$ cusps, respectively. Mazur [10, Theorem 1] shows that $\operatorname{Pic}(X)_{\operatorname{tors}}$ is a cyclic group of order n:=(p1)/(p1,12)n:=(p-1)/(p-1,12) generated by $P_{1}-P_{0}]$. It follows from [18, Proposition 5.4.1] that
$\langle P_{1}-P_{0},P_{1}-P_{0}\rangle=p\otimes\frac{12}{p-1}\in\mathbb{Q}^{\times}\otimes\mathbb{Q}/\mathbb{Z}.$
This can be computed by taking $f=(\eta(p\tau)/\eta(\tau))^{24/(p-1,12)}$ so that $\operatorname{div}(f)=n(P_{0}-P_{1})$. In particular, we have $\operatorname{Pic}(X)_{\operatorname{tors}}^{\operatorname{is}}=\{0\}$.
Example 2.6.
Let $p,q$ be distinct primes such that $p\equiv q\equiv 1\bmod 12$, and put
$a=\frac{(p-1)(q+1)}{24},\quad b=\frac{(p+1)(q-1)}{24},\quad c=\frac{(p-1)(q-1)}{24}.$
Let $X=X_{0}(pq)$ be the modular curve over $k=\mathbb{Q}$ and let $P_{0},P_{1},P_{2},P_{3}$ be the cusps of level $pq,1,p,q$, respectively. Ohta [12] shows $\operatorname{Pic}(X)_{\operatorname{tors}}$ is generated by $P_{i}-P_{0}\ (i=1,2,3)$ up to $2$-primary torsion. It follows from [18, Section 7] that the values of $\langle P_{i}-P_{0},P_{j}-P_{0}\rangle$ are given by $\otimes(1/2ab)$ of the following table.
[…]
In particular, $\operatorname{Pic}(X)_{\operatorname{tors}}^{\operatorname{is}}$ is a cyclic group of order $c$ generated by $P_{0}+P_{1}-P_{2}-P_{3}]$, up to $2$-primary torsion.
2.3. Functoriality
Let $Y$ be another geometrically irreducible smooth projective curve over $k$ and $\phi:X\to Y$ a finite morphism over $k$. We have the push-forward $\phi_{*}:\operatorname{Div}(X)\to\operatorname{Div}(Y)$ and the pull-back $\phi^{*}:\operatorname{Div}(Y)\to\operatorname{Div}(X)$ of divisors along $\phi$. Recall that they are given by
$\phi_{*}(\sum_{i}d_{i}P_{i})=\sum_{i}d_{i}[k(P_{i}):k(\phi(P_{i}))]\phi(P_{i}),$
$\phi^{*}(\sum_{j}e_{j}Q_{j})=\sum_{j}e_{j}\sum_{P\in\phi^{-1}(Q_{j})}{\operatorname{ram}}_{P}(\phi)P,$
where ${\operatorname{ram}}_{P}(\phi)$ denotes the ramification index of $\phi$ at $P$. We denote by $\langle\cdot,\cdot\rangle_{X}$ and $\langle\cdot,\cdot\rangle_{Y}$ the pairing (2.1) for $X$ and $Y$, respectively.
Lemma 2.7 (Projection formula).
For any $D\in\operatorname{Div}_{t}(X)$ and $E\in\operatorname{Div}^{0}(Y)$, we have
(2.6) $\langle\phi_{*}D,E\rangle_{X}=\langle D,\phi^{*}E\rangle_{Y}.$
Consequently, we have
$\phi^{*}(\operatorname{Pic}(Y)_{\operatorname{tors}}^{\operatorname{is}})\subset\operatorname{Pic}(X)_{\operatorname{tors}}^{\operatorname{is}},\qquad\phi_{*}(\operatorname{Pic}(X)_{\operatorname{tors}}^{\operatorname{is}})\subset\operatorname{Pic}(Y)_{\operatorname{tors}}^{\operatorname{is}}.$
Proof.
We may assume $|\phi_{*}D|\cap|E|=\emptyset$ and $|D|\cap|\phi^{*}E|=\emptyset$ by replacing $E$ by $E+\operatorname{div}(h)$ for a suitable $h\in k(Y)^{\times}$. Take $n\in\mathbb{Z}_{>0}$ and $f\in k(X)^{\times}$ such that $n[D]=0$ and $\operatorname{div}(f)=nD$. Since we have $\operatorname{div}(N_{k(X)/k(Y)}(f))=\phi_{*}\operatorname{div}(f)$, the left and right hand sides of (2.6) are rewritten as
$\prod_{j}(N_{k(X)/k(Y)}f)(Q_{j})^{e_{j}}\otimes\frac{1}{n},\qquad\prod_{j}\left(\prod_{P\in\phi^{-1}(Q_{j})}f(P)^{{\operatorname{ram}}_{P}(\phi)}\right)^{e_{j}}\otimes\frac{1}{n}.$
Now (2.6) follows from the formula $(N_{k(X)/k(Y)}f)(Q_{j})=\prod_{P\in\phi^{-1}(Q_{j})}f(P)^{{\operatorname{ram}}_{P}(\phi)}$. The last statement immediately follows from (2.6). We are done. ∎
Remark 2.8.
Let $k^{\prime}/k$ be a field extension, and $X^{\prime}$ the base change of $X$ to $k^{\prime}$. Then we have $\phi^{*}\langle D,E\rangle=\langle\phi^{*}D,\phi^{*}E\rangle^{\prime}$, where $\phi^{*}:k^{\times}\otimes\mathbb{Q}/\mathbb{Z}\to{k^{\prime}}^{\times}\otimes\mathbb{Q}/\mathbb{Z}$ and $\phi^{*}:\operatorname{Pic}(X)\to\operatorname{Pic}(X^{\prime})$ are the induced maps, and $\langle\cdot,\cdot\rangle^{\prime}$ denotes (2.1) for $X^{\prime}$. However, it can happen that $\phi^{*}(\operatorname{Pic}(X)_{\operatorname{tors}}^{\operatorname{is}})\not\subset\operatorname{Pic}(X^{\prime})_{\operatorname{tors}}^{\operatorname{is}}$ (when $\phi^{*}(\operatorname{Pic}(X)_{\operatorname{tors}})\subsetneq\operatorname{Pic}(X^{\prime})_{\operatorname{tors}}$).
3. Elliptic curves over $\mathbb{Q}$
Let $k$ be a field. Suppose that an elliptic curve $E:y^{2}+a_{1}xy+a_{3}y=x^{3}+a_{2}x^{2}+a_{4}x+a_{6}$ over $k$ has a $k$-rational point $P$ other than the point $O$ at infinity. By a suitable change of variables, we may assume that $P=(0,0)$, so the equation of the elliptic curve is of the form $y^{2}+a_{1}xy+a_{3}y=x^{3}+a_{2}x^{2}+a_{4}x$. We can show that $(0,0)$ is a $2$-torsion if and only if $a_{3}=0$. Thus, if $(0,0)$ is a $2$-torsion and $\operatorname{char}k\neq 2$, then we may make a change of variables $(x,y)\mapsto(x,y-a_{1}x/2)$ and assume that $E$ is of the form
(3.1) $E_{2,a,b}:y^{2}=x(x^{2}+ax+b),\qquad b(a^{2}-4b)\neq 0$
for some $a,b\in k$. If $(0,0)$ is not a $2$-torsion (without assumptions on $\operatorname{char}k$), then we have $a_{3}\neq 0$. We set $(x,y)\mapsto(x,y+a_{4}x/a_{3})$ and obtain an equation of the form $y^{2}+b_{1}xy+b_{3}y=x^{3}+b_{2}x^{2}$. We can show that $(0,0)$ is a $3$-torsion if and only if $b_{2}=0$. Thus, if $(0,0)$ is a $3$-torsion, then we may assume that the equation of the elliptic curve is of the form
(3.2) $E_{3,a,b}:y^{2}+axy+by=x^{3},\qquad b(a^{3}-27b)\neq 0$
for some $a,b\in k$. If $(0,0)$ is neither a $2$-torsion nor a $3$-torsion, then by setting $(x,y)\mapsto(u^{2}x,u^{3}y)$, where $u=b_{3}/b_{2}$, we find that the equation of the elliptic curve becomes
$E:y^{2}+(1+a)xy+by=x^{3}+bx^{2}$
for some $a,b\in k$. This is called the Tate normal form (with the starting rational point $P$) in the literature. By computing explicitly the coordinates of $m(0,0)$ for $m=1,2,\ldots$, one can determine the relation between $a$ and $b$ such that $(0,0)$ is of order $N$. As of now, the most extensive computation was due to Baaziz [1], who determined the relation for $N$ up to $51$. (Similar computation, but in a smaller range of $N$, can be found in several places, such as [13].) Such a relation can be taken as a defining equation for the modular curve $X_{1}(N)$. When $k$ is a subfield of $\mathbb{C}$, Baaziz also found expressions of $a$ and $b$ as modular functions in $\tau$, where $\tau\in\mathbb{H}$ is a point such that $(\mathbb{C}/(\mathbb{Z}\tau+\mathbb{Z}),1/N)\simeq(E,(0,0))$. The modular functions given in [1] were written in terms of the Weierstrass $\wp$-function. For our purpose, we will give alternative expressions using generalized Dedekind eta functions, which we recall in the next subsection.
Remark 3.1.
Note that the Tate normal form is unique in the sense that if
$E:y^{2}+(1+a)xy+by=x^{3}+bx^{2},\qquad E^{\prime}:y^{2}+(1+a^{\prime})xy+b^{\prime}y=x^{3}+b^{\prime}x^{2}$
are two elliptic curves in the Tate normal form such that $(E,(0,0))\simeq(E^{\prime},(0,0))$, then $a=a^{\prime}$ and $b=b^{\prime}$ (see Proposition 1.3 of [1]).
3.1. Generalized Dedekind eta functions
In this subsection, we recall the definition of generalized Dedekind eta functions and their properties [19].
Definition 3.2.
Let $N>1$ be an integer. For integers $g$ and $h$ not simultaneously congruent to 0 modulo $N$, the generalized Dedekind eta function $E_{g,h}^{(N)}$ is defined by
$E_{g,h}^{(N)}(\tau)=q^{B(g/N)/2}\prod_{m=1}^{\infty}(1-\zeta^{h}q^{m-1+g/N})(1-\zeta^{-h}q^{m-g/N}),$
where $q=e^{2\pi i\tau},\zeta=e^{2\pi i/N}$ and $B(x)=x^{2}-x+1/6$.
In the next lemma, we describe the properties of generalized Dedekind eta functions needed in our discussion.
Lemma 3.3 ([19, Theorem 1]).
Let $N>1$ be an integer and $g$ and $h$ be integers not simultaneously congruent to 0 modulo $N$. Then we have
(3.3) $E_{g+N,h}^{(N)}=E_{-g,-h}^{(N)}=-\zeta^{-h}E_{g,h}^{(N)},\qquad E_{g,h+N}^{(N)}=E_{g,h}^{(N)}.$
Also, for $\gamma=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in{\operatorname{SL}}(2,\mathbb{Z})$ with $c\geq 0$, we have
$E_{g,h}^{(N)}(\tau+b)=e^{\pi ibB(g/N)}E_{g,bg+h}^{(N)}(\tau),\qquad\text{if }c=0,$
and
(3.4) $E_{g,h}^{(N)}(\gamma\tau)=\epsilon(a,b,c,d)e^{\pi i\delta}E_{g^{\prime},h^{\prime}}^{(N)}(\tau),\qquad\text{if }c>0,$
where
$\epsilon(a,b,c,d)=\begin{cases}e^{\pi i(bd(1-c^{2})+c(a+d-3))/6},&\text{if }c\text{ is odd},\\ -ie^{\pi i(ac(1-d^{2})+d(b-c+3))/6},&\text{if }d\text{ is odd},\end{cases}$
$\delta=\frac{g^{2}ab+2ghbc+h^{2}cd}{N^{2}}-\frac{gb+h(d-1)}{N},$
and
$(g^{\prime}\ h^{\prime})=(g\ h)\begin{pmatrix}a&b\\ c&d\end{pmatrix}$.
Using Lemma 3.3, we easily obtain a criterion when a product of $E_{0,h}^{(N)}(\tau)$ is a modular function on $\Gamma_{1}(N)$.
Corollary 3.4.
Let $N>1$ be an integer. If $e_{1},\ldots,e_{N-1}$ are integers satisfying
$\sum_{h=1}^{N-1}e_{h}\equiv 0\bmod 12,\quad\sum_{h=1}^{N-1}he_{h}\equiv 0\bmod 2,\quad\sum_{h=1}^{N-1}h^{2}e_{h}\equiv 0\bmod 2N,$
then $\prod_{h=1}^{N-1}E_{0,h}^{(N)}(\tau)^{e_{h}}$ is a modular function on $\Gamma_{1}(N)$.
Proof.
Note that since $\left(\begin{smallmatrix}0&-1\\ N&0\end{smallmatrix}\right)$ normalizes $\Gamma_{1}(N)$, $\prod_{h=1}^{N-1}E_{0,h}^{(N)}(\tau)^{e_{h}}$ is modular on $\Gamma_{1}(N)$ if and only if $\prod_{h=1}^{N-1}E_{0,h}^{(N)}(-1/N\tau)^{e_{h}}$ is. By (3.3) and (3.4) applied to $\gamma=\left(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right)$, the latter is equal to $\prod_{h=1}^{N-1}E_{h,0}^{(N)}(N\tau)^{e_{h}}$, up to a constant. Then Corollary 3 of [19] shows that if $e_{1},\ldots,e_{N-1}$ satisfy the stated conditions, then $\prod_{h=1}^{N-1}E_{h,0}^{(N)}(N\tau)^{e_{h}}$ is a modular function on $\Gamma_{1}(N)$. Consequently, $\prod_{h=1}^{N-1}E_{0,h}^{(N)}(\tau)^{e_{h}}$ is a modular function on $\Gamma_{1}(N)$. This proves the corollary. ∎
Remark 3.5.
Note that by the Jacobi triple product identity, $E_{g,h}^{(N)}(\tau)$ is related to the Jacobi theta function $\vartheta_{1}(z|\tau)$ by
(3.5) $\vartheta_{1}\left(-\frac{g\tau+h}{N}\Big|\tau\right)=-ie^{-\pi ih/N}q^{-g^{2}/(2N^{2})}\eta(\tau)E_{g,h}^{(N)}(\tau),$
where $\eta(\tau)=q^{1/24}\prod_{n=1}^{\infty}(1-q^{n})$ is the classical Dedekind eta function. In particular, we have
(3.6) $\begin{split}\vartheta_{1}(h/N|\tau)&=-ie^{\pi ih/N}\eta(\tau)E_{0,-h}(\tau)=-ie^{\pi ih/N}\eta(\tau)(-e^{-2\pi ih/N})E_{0,h}^{(N)}(\tau)\\ &=ie^{-\pi ih/N}\eta(\tau)E_{0,h}^{(N)}(\tau),\end{split}$
which will be used frequently later on.
3.2. Hauptmodul
We now list our choice of Hauptmodul (generator of $\mathbb{C}(X_{1}(N))$) in the following lemma.
Lemma 3.6.
For $N\in\{2,\ldots,10,12\}$, the function $t(\tau)$ in the table below is a generator of the field of modular functions on $X_{1}(N)$ such that $t(\tau)$ takes value 0 and have a $\mathbb{Q}$-rational Fourier expansion at the cusp 0. In the table we also list the values of $t(\tau)$ at other cusps.
NN $t(\tau)$ values at other cusps 22 $\frac{\eta(\tau)^{24}}{64\eta(2\tau)^{24}}$ $t(\infty)=\infty$ 33 $\frac{\eta(\tau)^{12}}{27\eta(3\tau)^{12}}$ $t(\infty)=\infty$ 44 $\frac{1}{16}\frac{\eta(\tau)^{16}\eta(4\tau)^{8}}{\eta(2\tau)^{24}}$ $t(1/2)=\infty$ $t(\infty)=1/16$ 55 $\frac{E_{0,1}^{(5)}(\tau)^{5}}{E_{0,2}^{(5)}(\tau)^{5}}$ $t(1/2)=\infty$ $1-11t(a/5)-t(a/5)^{2}=0$ 66 $\frac{1}{9}\frac{\eta(\tau)^{8}\eta(6\tau)^{4}}{\eta(3\tau)^{8}\eta(2\tau)^{4}}$ $t(1/2)=1$, $t(1/3)=\infty$ $t(\infty)=1/9$ 77 $e^{4\pi i/7}\frac{E_{0,1}^{(7)}(\tau)^{3}}{E_{0,2}^{(7)}(\tau)^{2}E_{0,3}^{(7)}(\tau)}$ $t(1/2)=1$, $t(1/3)=\infty$ $1-8t(a/7)+5t(a/7)^{2}+t(a/7)^{3}=0$ 88 $i\frac{E_{0,1}^{(8)}(\tau)^{2}}{E_{0,3}^{(8)}(\tau)^{2}}$ $t(1/3)=\infty$, $t(1/2)=1$, t(1/4)=1t(1/4)=-1 $1-6t(a/8)+t(a/8)^{2}=0$ 99 $e^{4\pi i/9}\frac{E_{0,1}^{(9)}(\tau)^{2}}{E_{0,2}^{(9)}(\tau)E_{0,4}^{(9)}(\tau)}$ $t(1/2)=1$, $t(1/4)=\infty$, $1-t(a/3)+t(a/3)^{2}=0$ $1-6t(a/9)+3t(a/9)^{2}+t(a/9)^{3}=0$ $10$ $e^{4\pi i/10}\frac{E_{0,1}^{(10)}(\tau)E_{0,2}^{(10)}(\tau)}{E_{0,3}^{(10)}(\tau)E_{0,4}^{(10)}(\tau)}$ $t(1/3)=\infty$, $t(1/2)=1$, t(1/4)=1t(1/4)=-1 $1+t(a/5)-t(a/5)^{2}=0$, $1-4t(a/10)-t(a/10)^{2}=0$ $12$ $e^{4\pi i/12}\frac{E_{0,1}^{(12)}(\tau)}{E_{0,5}^{(12)}(\tau)}\phantom{\Bigg|}$ $t(1/5)=\infty$, $t(1/2)=1$, $1-t(a/3)+t(a/3)^{2}=0$ $1+t(a/4)^{2}=0$, t(1/6)=1t(1/6)=-1, $1-4t(a/12)+t(a/12)^{2}=0$
(Here in an expression such as $1-11t(a/5)-t(a/5)^{2}=0$, $a/5$ means any cusp represented by $a/5$ for some integer $a$ with $(a,5)=1$.)
Proof.
For $N=2,3,4,6$, we have $\pm\Gamma_{1}(N)=\Gamma_{0}(N)$, so $\mathbb{C}(X_{1}(N))=\mathbb{C}(X_{0}(N))$, where $\mathbb{C}(X(\Gamma))$ denotes the field of modular functions on a congruence subgroup $\Gamma$ of ${\operatorname{SL}}(2,\mathbb{Z})$. Generators of $\mathbb{C}(X_{0}(N))$ when $X_{0}(N)$ has genus 0 are well-known. For example, they are given in Table 3 of [3]. It is clear that $t(\infty)=\infty$ for $N=2,3$. Also, we may apply Propositions 3.2.1 and 3.2.8 of [9] to check that $t(\tau)=\eta(\tau)^{16}\eta(4\tau)^{8}/16\eta(2\tau)^{24}$ has only one simple zero at the cusp 0 and is nonvanishing elsewhere on $X_{0}(4)$. Then using the transformation formula $\eta(-1/\tau)=\sqrt{\tau/i}\eta(\tau)$, we see that the coefficients in the Fourier expansion of $t(\tau)$ at 0 are all rational numbers and the leading coefficient is $1$. The same thing can be said about the function $t(\tau)$ in the case $N=6$.
For the other cases, we shall use results from [19]. For the case $N=5$, applying the transformation formula in Lemma 3.3 with $\gamma=\left(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right)$, we find
$t(-1/5\tau)=-\frac{E_{0,1}^{(5)}(-1/5\tau)^{5}}{E_{0,2}^{(5)}(-1/5\tau)^{5}}=\frac{E_{1,0}^{(5)}(5\tau)^{5}}{E_{2,0}^{(5)}(5\tau)^{5}}.$
According to [19, Table 1], the function above generates $\mathbb{C}(X_{1}(5))$ and has a simple zero at the cusp $\infty$. Therefore, the function $t(\tau)$ itself generates $\mathbb{C}(X_{1}(5))$, has a simple zero at 0 and is nonvanishing elsewhere, and has a $\mathbb{Q}$-rational Fourier expansion with leading coefficient $1$ at the cusp 0. The other cases $N=7,8,9,10,12$ can be explained in the same way.
The values of $t(\tau)$ at cusps can be calculated using the transformation law for $E_{g,h}^{(N)}(\tau)$. Here as an example, we compute the value of $t(\tau)$ at $1/4$ in the case $N=10$. In the following computation, we omit the superscript $(10)$ from the notation $E_{g,h}^{(10)}(\tau)$.
Let $\gamma=\left(\begin{smallmatrix}1&0\\ 4&1\end{smallmatrix}\right)$. By Lemma 3.3, we have
$\begin{split}e^{4\pi i/10}\frac{E_{0,1}(\gamma\tau)E_{0,2}(\gamma\tau)}{E_{0,3}(\gamma\tau)E_{0,4}(\gamma\tau)}&=e^{4\pi i/10}\frac{e^{4\pi i/100}E_{4,1}(\tau)\cdot e^{16\pi i/100}E_{8,2}(\tau)}{e^{36\pi i/100}E_{12,3}(\tau)\cdot e^{64\pi i/100}E_{16,4}(\tau)}\\ &=e^{-4\pi i/10}\frac{E_{4,1}(\tau)E_{8,2}(\tau)}{(-e^{-6\pi i/10})E_{2,3}(\tau)\cdot(-e^{-8\pi i/10})E_{6,4}(\tau)}\\ &=-\frac{E_{4,1}(\tau)E_{8,2}(\tau)}{E_{2,3}(\tau)E_{6,4}(\tau)}=-1+\cdots.\end{split}$
Therefore, the value of the modular function $e^{4\pi i/10}E_{0,1}(\tau)E_{0,2}(\tau)/{E_{0,3}(\tau)E_{0,4}(\tau)}$ at the cusp $1/4$ is 1-1. We skip the details in other cases. ∎
We next give our choice of Hauptmodul of $X_{1}^{0}(N,2)$ when the modular curve has genus 0, i.e., when $N=2,4,6,8$.
Lemma 3.7.
For $N\in\{2,4,6,8\}$, the function $u(\tau)$ in the table below is a generator of the field of modular functions on $X_{1}^{0}(N,2)$ such that $u(\tau)$ takes value 0 and has a $\mathbb{Q}$-rational Fourier expansion at the cusp 0. In the table we also list the values of $u(\tau)$ at the other cusps.
NN $u(\tau)$ values at other cusps 22 $\frac{\eta(\tau/2)^{16}\eta(2\tau)^{8}}{\eta(\tau)^{24}}$ $u(\infty)=1$ $u(1)=\infty$ 44 $\frac{\eta(\tau/2)^{8}\eta(2\tau)^{4}}{\eta(\tau)^{12}}$ $u(1)=\infty$, u(1/2)=1u(1/2)=-1 $u(\infty)=1$ 66 $\frac{1}{3}\frac{\eta(\tau/2)^{4}\eta(3\tau)^{2}}{\eta(3\tau/2)^{4}\eta(\tau)^{2}}$ $u(1)=1$, u(1/2)=1u(1/2)=-1, u(1/3)=1/3u(1/3)=-1/3 $u(2/3)=\infty$, $u(\infty)=1/3$ 88 $e^{\pi i/4}\frac{E_{0,1}^{(8)}(\tau/2)}{E_{0,3}^{(8)}(\tau/2)}$ $u(1)=1$, $1+2u(a/4)-u(a/4)^{2}=0$, $1+u(a/2)^{2}=0$ u(1/3)=1u(1/3)=-1, $u(2/3)=\infty$, $1-2u(a/8)-u(a/8)^{2}=0$
Moreover, the function $u(\tau)$ is related to the generator $t(\tau)$ of $\mathbb{C}(X_{1}(N))$ given in Lemma 3.6 by
(3.7) $\begin{split}t=\begin{cases}4u/(1-u)^{2},&\text{when }N=2,\\ u/4(1+u)^{2},&\text{when }N=4,\\ u(1-u)/(1+3u),&\text{when }N=6,\\ u(1-u)/(1+u),&\text{when }N=8.\end{cases}\end{split}$
Proof.
We observe that when $N=2,4,6$, we have
$\pm\Gamma_{1}^{0}(N,2)=\begin{pmatrix}1&0\\ 0&2\end{pmatrix}^{-1}\Gamma_{0}(2N)\begin{pmatrix}1&0\\ 0&2\end{pmatrix}$.
We can check that $u(2\tau)$ generates $\mathbb{C}(X_{0}(2N))$, takes value 0 at the cusp 0, and has a $\mathbb{Q}$-rational Fourier expansion at 0. Therefore, $u(\tau)$ has the stated properties. The values of $u(\tau)$ at other cusps can be determined using the transformation law of $\eta(\tau)$.
We now consider the case $N=8$. For $\gamma=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\Gamma_{1}^{0}(8,2)$, we have
$\frac{1}{2}\gamma\tau=\begin{pmatrix}a&b/2\\ 2c&d\end{pmatrix}(\tau/2).$
Applying Lemma 3.3 with the matrix $\left(\begin{smallmatrix}a&b/2\\ 2c&d\end{smallmatrix}\right)$, we find that
$u(\gamma\tau)=e^{\pi i/4}\frac{E_{2c,d}^{(8)}(\tau/2)}{E_{6c,3d}^{(8)}(\tau/2)}=e^{\pi i/4}\frac{(-e^{-2\pi i/8})^{2c/8}E_{0,1}^{(8)}(\tau/2)}{(-e^{-6\pi i/8})^{2c/8}E_{0,3}^{(8)}(\tau/2)}=u(\tau).$
Thus, $u(\tau)$ is a modular function on $X_{1}^{0}(8,2)$. Using $q$-expansions, we can check that it is related to the Hauptmodul $t(\tau)$ for $X_{1}(8)$ specified in Lemma 3.6 by $t(\tau)=u(\tau)(1-u(\tau))/(1+u(\tau))$. Since $\mathbb{C}(X_{1}^{0}(8,2))$ is an extension of $\mathbb{C}(X_{1}(8))$ of degree $2$, $u(\tau)$ must be a generator of $\mathbb{C}(X_{1}^{0}(8,2))$. The values of $u(\tau)$ can be determined by the transformation law for generalized Dedekind eta functions. Finally, since $u(2\tau)^{2}=t(\tau)$, we see that $u(\tau)$ has $\mathbb{Q}$-rational Fourier expansions at the cusp 0.
The relation between $u(\tau)$ and $t(\tau)$ can be determined by using $q$-expansions or the covering $X_{1}^{0}(N,2)\to X_{1}(N)$. ∎
3.3. Parameterization of elliptic curves with $N$-torsion
In this subsection, we describe parameterizations of elliptic curves with prescribed rational torsions.
Lemma 3.8.
Let $N\geq 4$ be an integer. Let $E:y^{2}+(1+a)xy+by=x^{3}+bx^{2}$ be an elliptic curve over a subfield $k$ of $\mathbb{C}$ such that $(0,0)$ is of order $N$. Assume that $\tau$ is a point in $\mathbb{H}$ such that $(E,(0,0))\simeq_{\mathbb{C}}(\mathbb{C}/(\mathbb{Z}\tau+\mathbb{Z}),1/N)$. Then we have
(3.8) $\begin{split}a&=a(\tau)=\frac{\vartheta_{1}(1/N)^{4}\vartheta_{1}(4/N)}{\vartheta_{1}(2/N)^{5}}=\begin{cases}0,&\text{if }N=4,\\ \displaystyle e^{2\pi i/N}\frac{E_{0,1}^{(N)}(\tau)^{4}E_{0,4}^{(N)}(\tau)}{E_{0,2}^{(N)}(\tau)^{5}},&\text{if }N\geq 5,\end{cases}\\ b&=b(\tau)=\frac{\vartheta_{1}(1/N)^{5}\vartheta_{1}(3/N)^{3}}{\vartheta_{1}(2/N)^{8}}=e^{2\pi i/N}\frac{E_{0,1}^{(N)}(\tau)^{5}E_{0,3}^{(N)}(\tau)^{3}}{E_{0,2}^{(N)}(\tau)^{8}},\end{split}$
where we write the Jacobi theta function $\vartheta_{1}(z|\tau)$ simply as $\vartheta_{1}(z)$. The functions $a(\tau)$ and $b(\tau)$ are both modular functions on $\Gamma_{1}(N)$ with divisors supported on cusps.
Moreover, the elliptic curve $E$ is parameterized by
(3.9) $x(z)=-\frac{\vartheta_{1}(1/N)^{4}\vartheta_{1}(3/N)^{2}}{\vartheta_{1}(2/N)^{6}}\frac{\vartheta_{1}(z-1/N)\vartheta_{1}(z+1/N)}{\vartheta_{1}(z)^{2}}$
and
(3.10) $y(z)=\frac{\vartheta_{1}(1/N)^{7}\vartheta_{1}(3/N)^{3}}{\vartheta_{1}(2/N)^{10}}\frac{\vartheta_{1}(z-1/N)^{2}\vartheta_{1}(z+2/N)}{\vartheta_{1}(z)^{3}}.$
Proof.
According to Theorem 1.4 of [1], let
$a_{1}(z)=\frac{\wp^{\prime\prime}(z)}{\wp^{\prime}(z)},\qquad a_{2}(z)=3\wp(z)-\frac{1}{4}\frac{\wp^{\prime\prime}(z)^{2}}{\wp^{\prime}(z)^{2}},\qquad a_{3}(z)=\wp^{\prime}(z),$
where $\wp(z):=\wp(z|\tau)$ is the Weierstrass $\wp$-function, and $\widetilde{E}$ be the elliptic curve $y^{2}+a_{1}(1/N)xy+a_{3}(1/N)y=x^{3}+a_{2}(1/N)x^{2}$. Then $(\widetilde{E},(0,0))\simeq(\mathbb{C}/(\mathbb{Z}\tau+\mathbb{Z}),1/N)$ and the elliptic curve $\widetilde{E}$ is parameterized by
$x=\wp(z)-\wp(1/N),\qquad y=\frac{1}{2}(\wp^{\prime}(z)-a_{1}(1/N)x-a_{3}(1/N)).$
The elliptic curve is isomorphic to
$y^{2}+\frac{a_{1}(1/N)a_{2}(1/N)}{a_{3}(1/N)}xy+\frac{a_{2}(1/N)^{3}}{a_{3}(1/N)^{2}}y=x^{3}+\frac{a_{2}(1/N)^{3}}{a_{3}(1/N)^{2}}x^{2}$
through a simple change of variables with the parameterization of the new elliptic curve given by
(3.11) $x(z)=\frac{a_{2}(1/N)^{2}}{a_{3}(1/N)^{2}}(\wp(z)-\wp(1/N)),$
$y(z)=\frac{a_{2}(1/N)^{3}}{2a_{3}(1/N)^{3}}(\wp^{\prime}(z)-a_{1}(1/N)(\wp(z)-\wp(1/N))-a_{3}(1/N)).$
By the uniqueness of Tate normal forms (see Remark 3.1), we have
$a=\frac{a_{1}(1/N)a_{2}(1/N)-a_{3}(1/N)}{a_{3}(1/N)},\qquad b=\frac{a_{2}(1/N)^{3}}{a_{3}(1/N)^{2}}.$
Now Proposition 1.5 of [1] asserts that
$\operatorname{div}(a_{2})=\sum_{\begin{subarray}{c}3z=0\\ z\neq 0\end{subarray}}[z]-2\sum_{2z=0}[z],\qquad\operatorname{div}(a_{3})=\sum_{\begin{subarray}{c}2z=0\\ z\neq 0\end{subarray}}[z]-3[0],$
and
$\operatorname{div}(a_{1}a_{2}-a_{3})=\sum_{\begin{subarray}{c}4z=0\\ 2z\neq 0\end{subarray}}[z]-3\sum_{2z=0}[z].$
Therefore,
$a_{2}(z)=C_{1}(\tau)\frac{\vartheta_{1}(3z)}{\vartheta_{1}(z)\vartheta_{1}(2z)^{2}},\qquad a_{3}(z)=C_{2}(\tau)\frac{\vartheta_{1}(2z)}{\vartheta_{1}(z)^{4}},$
and
$a_{1}(z)a_{2}(z)-a_{3}(z)=C_{3}(\tau)\frac{\vartheta_{1}(4z)}{\vartheta_{1}(2z)^{4}}$
for some functions $C_{j}(\tau)$ of $\tau$. Considering the Laurent expansions $a_{2}(z)=\frac{3}{4}z^{-2}+\cdots$, $a_{3}(z)=-2z^{-3}+\cdots$, and $a_{1}(z)a_{2}(z)-a_{3}(z)=-\frac{1}{4}z^{-3}+\cdots$, we see that
(3.12) $a_{2}(z)=\frac{\vartheta_{1}^{\prime}(0)^{2}\vartheta_{1}(3z)}{\vartheta_{1}(z)\vartheta_{1}(2z)^{2}},\qquad a_{3}(z)=-\frac{\vartheta_{1}^{\prime}(0)^{3}\vartheta_{1}(2z)}{\vartheta_{1}(z)^{4}},$
and
$a_{1}(z)a_{2}(z)-a_{3}(z)=-\frac{\vartheta_{1}^{\prime}(0)^{3}\vartheta_{1}(4z)}{\vartheta_{1}(2z)^{4}}.$
It follows that
$a(\tau)=\frac{\vartheta_{1}(1/N)^{4}\vartheta_{1}(4/N)}{\vartheta_{1}(2/N)^{5}},\qquad b(\tau)=\frac{\vartheta_{1}(1/N)^{5}\vartheta_{1}(3/N)^{3}}{\vartheta_{1}(2/N)^{8}}.$
By (3.6), they can also be expressed as quotients of generalized Dedekind eta functions as claimed. Then applying Corollary 3.4, we easily see that $a(\tau)$ and $b(\tau)$ are modular functions on $\Gamma_{1}(N)$. Since $a(\tau)$ and $b(\tau)$ are products of generalized Dedekind eta functions, their zeros and poles must be at cusps. This proves the assertions about $a$ and $b$. We now find the expression for $x$ and $y$ in (3.11) in the Jacobi theta function.
Considering the zeros and poles of $\wp(z)-\wp(1/N)$, we find
$\wp(z)-\wp(1/N)=C_{4}(\tau)\frac{\vartheta_{1}(z-1/N)\vartheta_{1}(z+1/N)}{\vartheta_{1}(z)^{2}}$
for some function $C_{4}(\tau)$ of $\tau$. Since $\wp(z)-\wp(1/N)=z^{-2}+\cdots$ and $\vartheta_{1}(z)=\vartheta_{1}^{\prime}(0)z+\cdots$, we find that $C_{4}(\tau)=-\vartheta_{1}^{\prime}(0)^{2}/\vartheta_{1}(1/N)^{2}$. It follows that
$x(z)=\frac{a_{2}(1/N)^{2}}{a_{3}(1/N)^{2}}(\wp(z)-\wp(1/N))=-\frac{\vartheta_{1}(1/N)^{4}\vartheta_{1}(3/N)^{2}}{\vartheta_{1}(2/N)^{6}}\frac{\vartheta_{1}(z-1/N)\vartheta_{1}(z+1/N)}{\vartheta_{1}(z)^{2}}.$
For the elliptic function $y(z)$ in (3.11), we note that from the equation of the elliptic curve $E$, we see that $y(z)$ has a double zero at $1/N$. Since it has three zeros and a triple pole at 0, the other zero must be at 2/N-2/N. Thus,
$y(z)=C_{5}(\tau)\frac{\vartheta_{1}(z-1/N)^{2}\vartheta_{1}(z+2/N)}{\vartheta_{1}(z)^{3}}$
for some function $C_{5}(\tau)$ of $\tau$. Similar to the case of $x(z)$, by compairing the leading Laurent coefficients, we conclude the stated formula for $y(z)$. This proves the lemma. ∎
Proposition 3.9.
Let $N\in\{4,5,6,7,8,9,10,12\}$. Assume that $E$ is an elliptic curve over a subfield $k$ of $\mathbb{C}$ having a $k$-rational point $P$ of order $N$. Let $\tau$ be a point in $\mathbb{H}$ such that $(E,P)\simeq_{\mathbb{C}}(\mathbb{C}/(\mathbb{Z}\tau+\mathbb{Z}),1/N)$. Then $(E,P)\simeq_{k}(E_{N,t(\tau)},(0,0))$, where $t(\tau)$ is the Hauptmodul of $X_{1}(N)$ given in Lemma 3.6 and
(3.13) $E_{N,t}:y^{2}+(1+a(t))xy+b(t)y=x^{3}+b(t)x^{2}$
with $(a,b)$ given by
[…]
Proof.
Let $y^{2}+(1+a)xy+by=x^{3}+bx^{2}$ be the Tate normal form of $E$ with the starting point $P$. By Lemma 3.8 and the uniqueness of the Tate normal form, $a$ and $b$ are equal to $a(\tau)$ and $b(\tau)$ defined by (3.8). By the same lemma, they are modular functions on $\Gamma_{1}(N)$. Thus, they are rational functions in $t(\tau)$. For example, in the case $N=5$, by (3.3), we have
$a(\tau)=e^{2\pi i/5}\frac{E_{0,1}^{(5)}(\tau)^{4}E_{0,4}^{(5)}(\tau)}{E_{0,2}^{(5)}(\tau)^{5}}=-\frac{E_{0,1}^{(5)}(\tau)^{5}}{E_{0,2}^{(5)}(\tau)^{5}}=t(\tau),$
and similarly, $b(\tau)=t(\tau)$. In general, by compairing Fourier expansions, we see that the expressions of $a$ and $b$ in terms of $t$ are as stated. ∎
Remark 3.10.
The parameterization of elliptic curves with $N$-torsion obtained in the proposition is valid for any field $k$. In fact, such universal elliptic curves are already known in the literature (see, e.g., [5, 7]). The point of the proposition is to give explicit parameterization in terms of Hauptmodul of $X_{1}(N)$ when $k=\mathbb{C}$.
Note that the proof of the proposition itself does not explain why $E_{N,t}$ is defined over $k$ if and only if $t\in k$. This can be seen from the relation
$t=\begin{cases}b,&\text{for }N=4,5,\\ 1-b/a,&\text{for }N=6,\\ 1-b/a,&\text{for }N=7,\\ a/b-1,&\text{for }N=8,\\ (a-a^{2}-b)/(a-b),&\text{for }N=9,\\ (a-a^{2}-b)/a^{2},&\text{for }N=10,\\ -(a^{3}-ab+b^{2})/(a-b)^{2},&\text{for }N=12.\end{cases}$
Proposition 3.11.
Let $k$ be a subfield of $\mathbb{C}$. Let $N\in\{4,6,8\}$. Assume that $E$ is an elliptic curve over $k$ having $k$-rational points $P,Q$ such that $P$ is of order $N$, $Q$ is of order $2$ and such that $Q\notin\langle P\rangle$. Then there exists $\tau\in\mathbb{H}$ such that $(E,P,Q)\simeq_{k}(E^{\prime}_{N,u(\tau)},(0,0),Q^{\prime})$, where $u(\tau)$ is the Hauptmodul of $X_{1}^{0}(N,2)$ given in Lemma 3.7 and the equation of $E_{N,u}^{\prime}$ is given by
(3.14) $E^{\prime}_{N,u}:y^{2}+(1+a(u))xy+b(u)y=x^{3}+b(u)x^{2}$
with
$(a,b)=\begin{cases}\displaystyle\left(0,\frac{u}{4(1+u)^{2}}\right),&\text{if }N=4,\\ \displaystyle\left(\frac{u(1-u)}{1+3u},\frac{u(1-u)(1+u)^{2}}{(1+3u)^{2}}\right)&\text{if }N=6,\\ \displaystyle\left(\frac{u(1-u)(1+u^{2})}{(1+u)(1+2u-u^{2})},\frac{u(1-u)(1+u^{2})}{(1+2u-u^{2})^{2}}\right),&\text{if }N=8,\end{cases}$
and
(3.15) $Q^{\prime}=\begin{cases}\displaystyle\left(-\frac{1}{4(1+u)},\frac{1}{8(1+u)^{2}}\right),&\text{if }N=4,\\ \displaystyle\left(-\frac{(1-u)(1+u)^{2}}{4(1+3u)},\frac{(1-u)^{2}(1+u)^{3}}{8(1+3u)^{2}}\right),&\text{if }N=6,\\ \displaystyle\left(-\frac{(1-u)(1+u)(1+u^{2})}{4(1+2u-u^{2})},\frac{(1-u)^{2}(1+u)(1+u^{2})^{2}}{8(1+2u-u^{2})^{2}}\right),&\text{if }N=8.\end{cases}$
Proof.
Let $\tau$ be a point in $\mathbb{H}$ such that $(E,P)\simeq_{\mathbb{C}}(\mathbb{C}/(\mathbb{Z}\tau+\mathbb{Z}),1/N)$. The isomorphism yields $(E,P,Q)\simeq_{\mathbb{C}}(\mathbb{C}/(\mathbb{Z}\tau+\mathbb{Z}),1/N,\tau/2)$ or $(E,P,Q)\simeq_{\mathbb{C}}(\mathbb{C}/(\mathbb{Z}\tau+\mathbb{Z}),1/N,(\tau+1)/2)$. Changing $\tau$ to $\tau-1$ if necessary, we may assume that the first case occurs. Then by Lemma 3.8, we have $(E,P,Q)\simeq_{k}(E^{\prime},(0,0),Q^{\prime})$, where $E^{\prime}$ is the elliptic curve $y^{2}+a(\tau)xy+b(\tau)y=x^{3}+b(\tau)x^{2}$ with $a(\tau)$ and $b(\tau)$ given by (3.8) and $Q^{\prime}=(x(\tau/2),y(\tau/2))$ with $x(z)$ and $y(z)$ defined by (3.9) and (3.10), respectively. By the relation between $t(\tau)$ and $u(\tau)$ given in Lemma 3.7 and Proposition 3.9, we see that the equation $E^{\prime}$ is given by (3.14). Also, by (3.5), $x(\tau/2)$ and $y(\tau/2)$ can be written as
$\begin{split}x(\tau/2)&=-\frac{\vartheta_{1}(1/N)^{4}\vartheta_{1}(3/N)^{2}}{\vartheta_{1}(2/N)^{6}}\frac{\vartheta_{1}(\tau/2-1/N)\vartheta_{1}(\tau/2+1/N)}{\vartheta_{1}(\tau/2)^{2}}\\ &=-e^{2\pi i/N}\frac{E_{0,1}^{(N)}(\tau)^{4}E_{0,3}^{(N)}(\tau)^{2}}{E_{0,2}^{(N)}(\tau)^{6}}\frac{E_{-N/2,1}^{(N)}(\tau)E_{-N/2,-1}^{(N)}(\tau)}{E_{-N/2,0}^{(N)}(\tau)^{2}}\end{split}$
and
$\begin{split}y(\tau/2)&=\frac{\vartheta_{1}(1/N)^{7}\vartheta_{1}(3/N)^{3}}{\vartheta_{1}(2/N)^{10}}\frac{\vartheta_{1}(\tau/2-1/N)^{2}\vartheta_{1}(\tau/2+2/N)}{\vartheta_{1}(\tau/2)^{2}}\\ &=e^{4\pi i/N}\frac{E_{0,1}^{(N)}(\tau)^{7}E_{0,3}^{(N)}(\tau)^{3}}{E_{0,2}^{(N)}(\tau)^{10}}\frac{E_{-N/2,1}^{(N)}(\tau)^{2}E_{-N/2,-2}^{(N)}(\tau)}{E_{-N/2,0}^{(N)}(\tau)^{3}}.\end{split}$
Using (3.3), we simplify them to
$x(\tau/2)=-e^{2\pi i/N}\frac{E_{0,1}^{(N)}(\tau)^{4}E_{0,3}^{(N)}(\tau)^{2}}{E_{0,2}^{(N)}(\tau)^{6}}\frac{E_{N/2,1}^{(N)}(\tau)^{2}}{E_{N/2,0}^{(N)}(\tau)^{2}}$
and
$y(\tau/2)=e^{4\pi i/N}\frac{E_{0,1}^{(N)}(\tau)^{7}E_{0,3}^{(N)}(\tau)^{3}}{E_{0,2}^{(N)}(\tau)^{10}}\frac{E_{N/2,1}^{(N)}(\tau)^{2}E_{N/2,2}^{(N)}(\tau)}{E_{N/2,0}^{(N)}(\tau)^{3}}.$
Applying the transformation formula in Lemma 3.3, we may check that the two functions of $\tau$ are both modular functions on $\Gamma_{1}^{0}(N,2)$. Thus, they are both rational functions of $u(\tau)$. Compairing the Fourier expansions, we see that the expression of $Q^{\prime}$ in terms of $u$ is (3.15). ∎
Remark 3.12.
The parameterization is valid for any field $k$ with $\operatorname{char}k\neq 2$. To see this, we start from $(E,P)\stackrel{{\scriptstyle\phi}}{{\simeq}}_{k}(E_{N,t},(0,0))$, $t\in k$, where $E_{N,t}$ is given by (3.13) and $\phi$ is an isomorphism from $(E,P)$ to $(E_{N,t},(0,0))$. The elliptic curve $E_{N,t}$ has a full $2$-torsion subgroup if and only if the polynomial $x^{3}+bx^{2}+((1+a)x+b)^{2}/4$ splits completely over $k$. This implies that $t$ is of the form (3.7) for some $u\in k$. Since $\phi(Q)$ is a $2$-torsion of $E_{N,t}$ not in $\langle(0,0)\rangle$, we have
(3.16) $\phi(Q)=\begin{cases}\displaystyle\left(-\frac{1}{4(1+u)},\frac{1}{8(1+u)^{2}}\right),&\text{if }N=4,\\ \displaystyle\left(-\frac{(1-u)(1+u)^{2}}{4(1+3u)},\frac{(1-u)^{2}(1+u)^{3}}{8(1+3u)^{2}}\right),&\text{if }N=6,\\ \displaystyle\left(-\frac{(1-u)(1+u)(1+u^{2})}{4(1+2u-u^{2})},\frac{(1-u)^{2}(1+u)(1+u^{2})^{2}}{8(1+2u-u^{2})^{2}}\right),&\text{if }N=8\end{cases}$
(which is the coordinates of $Q^{\prime}$ in (3.15)), or
(3.17) $\phi(Q)=\begin{cases}\displaystyle\left(-\frac{u}{4(1+u)},\frac{u^{2}}{8(1+u)^{2}}\right),&\text{if }N=4,\\ \displaystyle\left(-\frac{u(1+u)^{2}}{(1+3u)^{2}},\frac{u^{2}(1+u)^{3}}{(1+3u)^{3}}\right),&\text{if }N=6,\\ \displaystyle\left(-\frac{u(1+u^{2})}{(1+u)^{2}(1+2u-u^{2})},\frac{u^{2}(1+u^{2})^{2}}{(1+u)^{3}(1+2u-u^{2})^{3}}\right),&\text{if }N=8.\end{cases}$
Observe that $t$ is invariant under the substitution
$u\longmapsto\begin{cases}1/u,&\text{if }N=4,\\ (1-u)/(1+3u),&\text{if }N=6,\\ (1-u)/(1+u),&\text{if }N=8,\end{cases}$
while the coordinates in (3.17) is changed to those in (3.16) under the substitution. (In the case $k=\mathbb{C}$, this corresponds to the automorphism $\tau\mapsto\tau+1$ on $X_{1}^{0}(N,2)$.) This implies that $(E,P,Q)\simeq_{k}(E^{\prime}_{N,u},(0,0),Q^{\prime})$ for some $u\in k$.
4. Computation of the pairings
In this section, we conpute the pairing $\langle P,Q\rangle$ for $\mathbb{Q}$-rational torsion points $P,Q$ on an elliptic curve $E$ over $\mathbb{Q}$. (By our convention, $\langle P,Q\rangle$ here means what is denoted by $\langle P-O,Q-O\rangle$ in (2.1), where $O\in E(k)$ is the identity element.) The results are described in terms of universal polynomials $f_{N}(t),g_{N}^{(\nu)}(u)$. They will be used to construct new modular curves $X_{1}(N)_{M}^{\pm}$ and $X_{1}^{0}(N,2)_{\underline{M}}^{\pm}$ in the next section.
4.1. Cyclic case
In the following discussion, for a nonzero rational function $F$ on an elliptic curve over a field $k$ and a point $Q\in E(k)$, we let ${\operatorname{LC}}_{Q,\pi}(F)$ denote the leading coefficient of $F$ at $Q$ with respect to a uniformizer $\pi$ at $Q$.
Proposition 4.1.
Let $N\geq 4$ be an integer. Let $E:y^{2}+(1+a)xy+by=x^{3}+bx^{2}$ be an elliptic curve over a subfield $k$ of $\mathbb{C}$ such that $P:=(0,0)$ is of order $N$. Assume that $\tau$ is a point in $\mathbb{H}$ such that $(E,P)\simeq(\mathbb{C}/(\mathbb{Z}\tau+\mathbb{Z}),1/N)$. Let $F(x,y)$ be a rational function on $E$ such that $\operatorname{div}(F)=N(P-O)$. Choose $x$ and $y/x$ to be the local parameters for the points $P$ and $O$, respectively. Then
$\frac{{\operatorname{LC}}_{P,x}(F)}{{\operatorname{LC}}_{O,y/x}(F)}=\left(\frac{\vartheta_{1}(2/N)^{2}}{\vartheta_{1}(1/N)\vartheta_{1}(3/N)}\right)^{N}=\left(\frac{E_{0,2}^{(N)}(\tau)^{2}}{E_{0,1}^{(N)}(\tau)E_{0,3}^{(N)}(\tau)}\right)^{N}.$
Moreover, the last expression is a modular function on $\Gamma_{1}(N)$.
Proof.
Let $x(z)$ and $y(z)$ be defined by (3.9) and (3.10), respectively. Then
$F(x(z),y(z))=C\left(\frac{\vartheta_{1}(z-1/N)}{\vartheta_{1}(z)}\right)^{N}$
for some constant $C$. We have
$\begin{split}{\operatorname{LC}}_{P,x}(F)&=C\lim_{z\to 1/N}x(z)^{-N}\left(\frac{\vartheta_{1}(z-1/N)}{\vartheta_{1}(z)}\right)^{N}\\ &=C\lim_{z\to 1/N}\left(-\frac{\vartheta_{1}(2/N)^{6}\vartheta_{1}(z)^{2}}{\vartheta_{1}(1/N)^{4}\vartheta_{1}(3/N)^{2}\vartheta_{1}(z-1/N)\vartheta_{1}(z+1/N)}\frac{\vartheta_{1}(z-1/N)}{\vartheta_{1}(z)}\right)^{N}\\ &=C\left(-\frac{\vartheta_{1}(2/N)^{5}}{\vartheta_{1}(1/N)^{3}\vartheta_{1}(3/N)^{2}}\right)^{N}.\end{split}$
Similarly, we compute that
$\begin{split}{\operatorname{LC}}_{O,y/x}(F)&=C\lim_{z\to 0}\left(\frac{y(z)}{x(z)}\right)^{-N}\left(\frac{\vartheta_{1}(z-1/N)}{\vartheta_{1}(z)}\right)^{N}\\ &=C\lim_{z\to 0}\left(-\frac{\vartheta_{1}(2/N)^{4}\vartheta_{1}(z)\vartheta_{1}(z+1/N)}{\vartheta_{1}(1/N)^{3}\vartheta_{1}(3/N)\vartheta_{1}(z-1/N)\vartheta_{1}(z+2/N)}\frac{\vartheta_{1}(z-1/N)}{\vartheta_{1}(z)}\right)^{N}\\ &=C\left(-\frac{\vartheta_{1}(2/N)^{3}}{\vartheta_{1}(1/N)^{2}\vartheta_{1}(3/N)}\right)^{N}.\end{split}$
It follows that
$\frac{{\operatorname{LC}}_{P,x}(F)}{{\operatorname{LC}}_{O,y/x}(F)}=\left(\frac{\vartheta_{1}(2/N)^{2}}{\vartheta_{1}(1/N)\vartheta_{1}(3/N)}\right)^{N}.$
By (3.5), it can also be written as
$\left(\frac{E_{0,2}^{(N)}(\tau)^{2}}{E_{0,1}^{(N)}(\tau)E_{0,3}^{(N)}(\tau)}\right)^{N}.$
Applying Corollary 3.4, we see that the function above is a modular function on $\Gamma_{1}(N)$. This completes the proof of the proposition. ∎
We now compute the pairing $\langle P,P\rangle$ for an $N$-torsion rational point $P$. We treat the cases $N=2$ and $3$ separately in §4.3. For other $N$, the pairing is given in the next proposition.
Proposition 4.2.
Let $k$ be a subfield of $\mathbb{C}$. Assume that $N\in\{4,5,6,7,8,9,10,12\}$. Let $E_{N,t}$, $t\in k$, be the elliptic curve defined by (3.13) and put $P:=(0,0)\in E_{N,t}(k)_{\operatorname{tors}}$. Then
$\langle P,P\rangle=f_{N}(t)\otimes(1/N),$
where $f_{N}(t)$ is given by
[…]
Proof.
By Proposition 4.1 above, for the elliptic curve $E_{N,t(\tau)}$, we have
$\langle P,P\rangle=\left(\frac{E_{0,1}^{(N)}(\tau)E_{0,3}^{(N)}(\tau)}{E_{0,2}^{(N)}(\tau)^{2}}\right)^{N}\otimes(1/N).$
Let $f(\tau)$ be the function in the expression above. Since $f(\tau)$ is a modular function on $\Gamma_{1}(N)$, it is a rational function of $t(\tau)$. By computing the Fourier expansions, we find that $f(\tau)=f_{N}(t(\tau))$ for $N=4,5,6,7,9$, and
$f(\tau)=\begin{cases}\displaystyle\frac{t(1-t)^{2}}{(1+t)^{4}},&\text{when }N=8,\\ \displaystyle\frac{t(1-t)^{2}}{(1+t)^{2}(1+t-t^{2})^{5}},&\text{when }N=10,\\ \displaystyle\frac{t(1-t)^{2}(1-t+t^{2})^{3}(1+t^{2})^{4}}{(1+t)^{6}},&\text{when }N=12.\end{cases}$
This yields the claimed formula for the pairing $\langle P,P\rangle$. ∎
4.2. Non-cyclic case
We now compute the pairing $\langle P,P\rangle,\langle Q,Q\rangle,\langle P,Q\rangle$ for $P\in E(\mathbb{Q})[N]$ and $Q\in E(\mathbb{Q})[2]$ such that $Q\not\in\langle P\rangle$. Again, we treat the case $N=2$ separately in §4.3. The cases $N=4,6,8$ are treated in the following proposition.
Proposition 4.3.
Let $k$ be a subfield of $\mathbb{C}$ and $N\in\{4,6,8\}$. Let $E$ be an elliptic curve defined by (3.14), $P:=(0,0)$ and $Q$ the 22-torsion point given in (3.15). Then we have
$\langle P,P\rangle=g_{N}^{(1)}(u)\otimes\frac{1}{N},\quad\langle Q,Q\rangle=g_{N}^{(2)}(u)\otimes\frac{1}{2},\quad\langle P,Q\rangle=g_{N}^{(3)}(u)\otimes\frac{1}{2},$
where $g_{N}^{(\nu)}(u)$ is given by
[…]
Proof.
The formula for $\langle P,P\rangle$ follows from Proposition 4.2 and (3.7). We next compute $\langle Q,Q\rangle$. Let $(x_{0},y_{0})$ be the coordinates of $Q$ given in (3.15). We have $\operatorname{div}(x-x_{0})=2(Q-O)$. Choose $y^{\prime}=y+(1+a)x/2+b/2$ and $y/x$ to be uniformizers at $Q$ and $O$, respectively. Then
$\langle Q,Q\rangle=\frac{{\operatorname{LC}}_{Q,y^{\prime}}(x-x_{0})}{{\operatorname{LC}}_{O,y/x}(x-x_{0})}\otimes(1/2).$
Now we have
$\frac{x-x_{0}}{(y^{\prime})^{2}}=\left(x^{2}+\left(x_{0}+\frac{(1+a)^{2}}{4}+b\right)x-\frac{b^{2}}{4x_{0}}\right)^{-1}$
and consequently,
${\operatorname{LC}}_{Q,y^{\prime}}(x-x_{0})=\left(2x_{0}^{2}+\frac{(1+a)^{2}x_{0}}{4}+bx_{0}-\frac{b^{2}}{4x_{0}}\right)^{-1}.$
Plugging in the actual values of $x_{0}$, $a$, and $b$, we find
${\operatorname{LC}}_{Q,y^{\prime}}(x-x_{0})=\begin{cases}\displaystyle\frac{1-u}{16(1+u)^{3}},&\text{if }N=4,\\ \displaystyle\frac{(1-u)^{3}(1+u)^{3}(1-3u)}{16(1+3u)^{3}},&\text{if }N=6,\\ \displaystyle\frac{(1-2u-u^{2})(1+u^{2})^{2}(1-u)^{4}}{16(1+2u-u^{2})^{3}},&\text{if }N=8,\end{cases}$
On the other hand, it is clear that ${\operatorname{LC}}_{O,y/x}(x-x_{0})=1$. Then the formula for $\langle Q,Q\rangle$ follows.
To compute $\langle P,Q\rangle$, we choose $x$ and $y/x$ to be the local parameters at $P$ and $O$, respectively. Then ${\operatorname{LC}}_{P,x}(x-x_{0})$ is simply $x_{0}$ and hence $\langle P,Q\rangle=(-x_{0})\otimes(1/2)$, which yields the claimed formula for $\langle P,Q\rangle$. ∎
4.3. Torsion points of order two and three
We need to treat these cases with special care, since $X_{1}(2),X_{1}(3),X_{1}^{0}(2,2)$ are not fine moduli. The following lemma is valid over a general field $k$ (subject to a condition on the characteristic).
Lemma 4.4.
  1. (1)
    Assume that the characteristic of $k$ is not $2$. If $E$ is an elliptic curve over $k$ with a $k$-rational point $P$ of order two, then $(E,P)$ is isomorphic to $(E_{2,t,a},(0,0))$ for some $a,t\in k^{\times}$, where
    (4.1) $E_{2,t,a}:\begin{cases}\displaystyle y^{2}=x\left(x^{2}+ax+\frac{a^{2}t}{4(t+1)}\right),&\text{if }t\neq-1,\\ y^{2}=x(x^{2}+a),&\text{if }t=-1.\end{cases}$
    Moreover, two elliptic curves $(E_{2,t,a},(0,0))$ and $(E_{2,t^{\prime},a^{\prime}},(0,0))$ are isomorphic over $k$ if and only if $t=t^{\prime}\neq-1$ and $a/a^{\prime}$ is a square in $k^{\times}$ or $t=t^{\prime}=-1$ and $a/a^{\prime}\in(k^{\times})^{4}$ .
  2. (2)
    Assume that the characteristic of $k$ is not $3$. If $E$ is an elliptic curve over $k$ with a $k$-rational point $P$ of order $3$, then $(E,P)$ is isomorphic to $(E_{3,t},(0,0))$ for some $t\neq-1$ in $k^{\times}$, where
    (4.2) $E_{3,t}:y^{2}+xy+\frac{t}{27(t+1)}y=x^{3}$
    or isomorphic to $(E_{3,-1,a},(0,0))$ for some $a\in k^{\times}$, where
    (4.3) $E_{3,-1,a}:y^{2}+ay=x^{3}.$
    Two elliptic curves $(E_{3,t},(0,0))$ and $(E_{3,t^{\prime}},(0,0))$ with $t,t^{\prime}\neq-1$ are isomorphic if and only if $t=t^{\prime}$. Also, $(E_{3,-1,a},(0,0))\simeq(E_{3,-1,a^{\prime}},(0,0))$ if and only if $a/a^{\prime}\in(k^{\times})^{3}$.
Proof.
Consider first the case $N=2$. We start with the equation (3.1) and introduce a new parameter $t=4b/(a^{2}-4b)$ (so $b=a^{2}t/4(t+1)$). When $a\neq 0$, we have $t\neq-1$ and the equation (3.1) in the new parameters is $y^{2}=x(x^{2}+ax+a^{2}t/4(t+1))$. When $a=0$, we change the notation $b$ in (3.1) to a-a and get $y^{2}=x(x^{2}-a)$. Now it is easy to see that an isomorphism $\phi:(E_{2,t,a},(0,0))\to(E_{2,t^{\prime},a^{\prime}},(0,0))$ must be of the form $\phi(x,y)=(u^{2}x,u^{3}y)$ (c.f. [15, Proposition 3.1]). From this, we immediately get the condition for $E_{2,t,a}$ and $E_{2,t^{\prime},a^{\prime}}$ to be isomorphic.
For the case $N=3$, we similarly let $t=27b/(a^{3}-27b)$ (so $b=a^{3}t/27(t+1)$). Then (3.2) becomes
$y^{2}+axy+\frac{a^{3}t}{27(t+1)}y=x^{3}.$
If $a\neq 0$, i.e., if $t\neq-1$, we make a change of variables $(x,y)\mapsto(a^{2}x,a^{3}y)$. Then the equation reduces to $E_{3,t}$. If $a=0$, i.e., if t=1t=-1, then we change $b$ to $a$ and obtain $E_{3,-1,a}$. It is straightforward to verify the conditions for elliptic curves to be isomorphic. ∎
Remark 4.5.
When $k=\mathbb{C}$ and $\tau\in\mathbb{C}$, one can show $(E_{2,t(\tau),a},(0,0))\simeq(\mathbb{C}/(\mathbb{Z}\tau+\mathbb{Z}),1/2)$ and $(E_{3,t(\tau)},(0,0))\simeq(\mathbb{C}/(\mathbb{Z}\tau+\mathbb{Z}),1/3)$ if $t(\tau)\neq-1$ and $(E_{3,-1,a},(0,0))\simeq(\mathbb{C}/(\mathbb{Z}\frac{1+\sqrt{-3}}{2}+\mathbb{Z}),1/3)$, where $t(\tau)$ is given as in Lemma 3.6. We omit the proof, as we will not use them.
We compute the pairing in the following lemma.
Lemma 4.6.
Let $k$ be a subfield of $\mathbb{C}$.
  1. (1)
    Let $E_{2,t,a}$ be an elliptic curve (4.1) for some $a,t\in k^{\times}$, and put $P:=(0,0)$. Then
    $\langle P,P\rangle=\begin{cases}t(1+t)\otimes\frac{1}{2},&\text{if }t\neq-1,\\ a\otimes\frac{1}{2},&\text{if }t=-1.\end{cases}$
  2. (2)
    Let $t\in k^{\times}$, and if t=1t=-1 take also $a\in k^{\times}$. Let $E$ be an elliptic curve given by either $E_{3,t}$ from (4.2) or $E_{3,-1,a}$ from (4.3), and put $P:=(0,0)$. Then
[…]
Proof.
(1) The divisor of the function $x$ on $E_{2,t,a}$ is 2(PO)2(P-O). We choose $y$ to be a uniformizer at $P$ and $y/x$ to be a uniformizer at $O$. Then
${\operatorname{lc}}_{P}(x,2)=\frac{x}{y^{2}}\Big|_{P}\otimes\frac{1}{2}=\begin{cases}\displaystyle\frac{4(1+t)}{a^{2}t}\otimes\frac{1}{2},&\text{if }t\neq-1,\\ \displaystyle\frac{1}{a}\otimes\frac{1}{2},&\text{if }t=-1,\end{cases}$
${\operatorname{lc}}_{O}(x,2)=\frac{x}{(y/x)^{2}}\Big|_{O}\otimes\frac{1}{2}=0.$
Therefore, we have $\langle P,P\rangle=t(1+t)\otimes(1/2)$ when $t\neq-1$ and $\langle P,P\rangle=a\otimes(1/2)$ when t=1t=-1.
(2) The divisor of the function $y$ on $E$ is 3(PO)3(P-O). We choose $x$ to be a uniformizer at $P$ and $x/y$ to be a uniformizer at $O$. Then
${\operatorname{lc}}_{P}(y,3)=\frac{y}{x^{3}}\Big|_{P}\otimes\frac{1}{3}=\begin{cases}\displaystyle\frac{27(1+t)}{t}\otimes\frac{1}{3},&\text{if }t\neq-1,\\ \displaystyle\frac{1}{a}\otimes\frac{1}{3},&\text{if }t=-1,\end{cases}$
${\operatorname{lc}}_{O}(y,3)=\frac{y}{(y/x)^{3}}\Big|_{O}\otimes\frac{1}{3}=0.$
It follows that $\langle P,P\rangle=t(1+t)^{2}\otimes(1/3)$ when $t\neq-1$ and $\langle P,P\rangle=a\otimes(1/3)$ when t=1t=-1. ∎
Next, we consider elliptic curves with full $2$-torsion points. It is clear that, if $E$ is an elliptic curve over a field $k$ with $\operatorname{char}k\not=2$ having two distinct $k$-rational 22-torsion points $P$ and $Q$, then $(E,P,Q)\simeq(E^{\prime}_{2,u,a},(0,0),(a,0))$ for some $u\neq 0,1\in k$ and $a\in k^{\times}$, where
(4.4) $E_{2,u,a}^{\prime}:y^{2}=x(x-a)(x-au).$
Lemma 4.7.
Let $E=E_{2,u,a}^{\prime}$ be an elliptic curve from (4.4) over a subfield $k$ of $\mathbb{C}$, and put $P:=(0,0),Q:=(a,0)$. Then we have
$\langle P,P\rangle=u\otimes\frac{1}{2},\quad\langle Q,Q\rangle=(1-u)\otimes\frac{1}{2},\quad\langle P,Q\rangle=a\otimes\frac{1}{2}.$
Proof.
We have $\operatorname{div}x=2(P-O)$ and $\operatorname{div}(x-a)=2(Q-O)$. Choose $y$, $y$, $x/y$ to be uniformizers at $P$, $Q$, and $O$, respectively. We have
${\operatorname{lc}}_{P}(x,2)=\frac{x}{y^{2}}\Big|_{P}\otimes\frac{1}{2}=\frac{1}{(x-a)(x-au)}\Big|_{P}\otimes\frac{1}{2}=u\otimes\frac{1}{2},$
${\operatorname{lc}}_{Q}(x-a,2)=\frac{x-a}{y^{2}}\Big|_{Q}\otimes\frac{1}{2}=\frac{1}{x(x-au)}\Big|_{Q}\otimes\frac{1}{2}=(1-u)\otimes\frac{1}{2},$
${\operatorname{lc}}_{P}(x-a,2)=(-a)\otimes\frac{1}{2},$
${\operatorname{lc}}_{O}(x,2)=\frac{x}{(y/x)^{2}}\Big|_{O}\otimes\frac{1}{2}=0,$
${\operatorname{lc}}_{O}(x-a,2)=\frac{x-a}{(y/x)^{2}}\Big|_{O}\otimes\frac{1}{2}=0.$
Then the formulas for the pairing follows. ∎
Remark 4.8.
When $k=\mathbb{C}$, one can show $(\mathbb{C}/\Lambda_{\tau},1/2,\tau/2)\simeq(E_{2,u(\tau),a}^{\prime},(0,0),(a,0))$, where $\tau\in\mathbb{C}$ and $u(\tau)$ is the uniformizer of $X_{1}^{0}(2,2)=X(2)$ given in Lemma 3.7. We will not use this fact.
5. Modular curves $X_{1}(N)_{M}^{\pm}$ and $X_{1}^{0}(N,2)_{\underline{M}}^{\pm}$
Using the universal polynomials from Propositions 4.2 and 4.3, we are now able to construct new modular curves that encode information on the values of the pairing.
5.1. Modular curve $X_{1}(N)_{M}^{\pm}$
Definition 5.1.
Assume $N\in\{4,5,6,7,8,9,10,12\}$, let $M$ be a positive divisor of $N$, and $\epsilon\in\{+,-\}$. Let $f_{N}(t)$ be the function from Proposition 4.2. We define $X_{1}(N)_{M}^{\epsilon}$ to be the smooth projective model over $\mathbb{Q}$ of $\epsilon s^{M}=f_{N}(t)$. In other words, $X_{1}(N)_{M}^{\epsilon}$ is the normalization of the fiber product
$X_{1}^{\prime}(N)_{M}^{\epsilon}:=X_{1}(N)\times_{\mathbb{P}^{1}_{z}}\mathbb{P}^{1}_{s},$
where $X_{1}(N)\to\mathbb{P}^{1}_{z}$ is the composition of the isomorphism $X_{1}(N)\cong\mathbb{P}^{1}_{t}$ from Lemma 3.6 followed by $\mathbb{P}^{1}_{t}\to\mathbb{P}^{1}_{z},\ z=f_{N}(t)$, and $\mathbb{P}^{1}_{s}\to\mathbb{P}^{1}_{z}$ is defined by $z=\epsilon s^{M}$. We write $Y_{1}(N)_{M}^{\epsilon}\subset X_{1}(N)_{M}^{\epsilon}$ for the inverse image of $Y_{1}(N)$, and put $X_{1}(N)_{M}^{\pm}:=X_{1}(N)_{M}^{+}\sqcup X_{1}(N)_{M}^{-}$, $Y_{1}(N)_{M}^{\pm}:=Y_{1}(N)_{M}^{+}\sqcup Y_{1}(N)_{M}^{-}$. Here, as usual, we denote the complement of cusps in $X_{1}(N)$ by $Y_{1}(N)$.
The motivation for the definition will become clear in Corollary 5.5 below.
Remark 5.2.
  1. (1)
    If $k$ is a field of characteristic zero containing an element $\zeta$ such that $\zeta^{M}=-1$, there is an isomorphism $(X_{1}(N)_{M}^{+})_{k}\cong(X_{1}(N)_{M}^{-})_{k}$ given by $s\mapsto\zeta s$, where $(-)_{k}$ denotes the base change by $k/\mathbb{Q}$. In particular, if $M$ is odd we have $X_{1}(N)_{M}^{+}\cong X_{1}(N)_{M}^{-}$ over $\mathbb{Q}$ (by taking $\zeta=-1$).
  2. (2)
    Observe that $t=0$ is a simple zero of $f_{N}(t)$ for any $N$. It follows that $X_{1}^{\prime}(N)_{M}^{\epsilon}$ is non-singular at $(s,t)=(0,0)\in X_{1}^{\prime}(N)_{M}^{\epsilon}(\mathbb{Q})$, and hence there is a unique $\mathbb{Q}$-rational point on $X_{1}(N)_{M}^{\epsilon}$ above it.
  3. (3)
    For a positive divisor $M^{\prime}$ of $M$, there is a finite morphism
    $X_{1}(N)_{M}^{\epsilon}\to X_{1}(N)_{M^{\prime}}^{\epsilon}$
    induced by $s\mapsto s^{M/M^{\prime}}$. In particular, by taking $M^{\prime}=1$ we get
    $X_{1}(N)_{M}^{\epsilon}\to X_{1}(N)_{1}^{\epsilon}=X_{1}(N).$
The normalization process is unnecessary for $Y_{1}(N)_{N}^{\epsilon}$.
Lemma 5.3.
The canonical map $Y_{1}(N)_{M}^{\epsilon}\to X_{1}^{\prime}(N)_{M}^{\epsilon}$ is an isomorphism onto its image.
Proof.
Observe that the branch locus of $\mathbb{P}^{1}_{s}\to\mathbb{P}^{1}_{z},\ z=\epsilon s^{M}$ is contained in $\{0,\infty\}\subset\mathbb{P}^{1}_{z}$ (equal if $M>1$), and that all points in $f_{N}^{-1}(\{0,\infty\})\subset\mathbb{P}^{1}_{t}=X_{1}(N)$ are cusps by Lemma 3.6. We conclude $X_{1}^{\prime}(N)_{M}^{\epsilon}$ is smooth over $\mathbb{Q}$ outside cusps, whence the lemma. ∎
Recall that we have specified an isomorphism $t:X_{1}(N)\cong\mathbb{P}^{1}$ in Lemma 3.6. By the moduli interpretation, there is a bijection between $Y_{1}(N)(\mathbb{Q})$ and the set of isomorphism classes of pairs $(E,P)$ of an elliptic curve $E$ over $\mathbb{Q}$ and a $\mathbb{Q}$-rational point $P$ of order $N$. By abuse of notation, we write $(E,P)\in Y_{1}(N)(\mathbb{Q})$ in this situation, and write $t(E,P)\in\mathbb{Q}^{\times}\subset\mathbb{P}^{1}(\mathbb{Q})$ for the point corresponding to $(E,P)\in Y_{1}(N)(\mathbb{Q})$. (We have $t(E,P)\not=0,\infty$ since $t=0,\infty$ correspond to cusps on $X_{1}(N)$ by Lemma 3.6.)
Lemma 5.4.
Let $N,M,\epsilon$ be as in Definition 5.1.
  1. (1)
    There is a bijection between $Y_{1}(N)_{M}^{\epsilon}(\mathbb{Q})$ and the set of triples $(E,P,s)$ where $(E,P)\in Y_{1}(N)(\mathbb{Q})$ and $s\in\mathbb{Q}^{\times}$ such that $\epsilon s^{M}=f_{N}(t(E,P))$.
  2. (2)
    Given $(E,P)\in Y_{1}(N)(\mathbb{Q})$, we have
    $\langle P,(N/M)P\rangle=0\Leftrightarrow(E,P)\in\operatorname{Im}(Y_{1}(N)_{M}^{\pm}(\mathbb{Q})\to Y_{1}(N)(\mathbb{Q})).$
    If $M$ is odd, one may replace $Y_{1}(N)_{M}^{\pm}(\mathbb{Q})$ by $Y_{1}(N)_{M}^{+}(\mathbb{Q})$.
  3. (3)
    Suppose $M^{\prime}|M$. Then $X_{1}(N)_{M}^{\epsilon}\cong\mathbb{P}^{1}$ implies $X_{1}(N)_{M^{\prime}}^{\epsilon}\cong\mathbb{P}^{1}$, and $Y_{1}(N)_{M^{\prime}}^{\epsilon}(\mathbb{Q})=\emptyset$ implies $Y_{1}(N)_{M}^{\epsilon}(\mathbb{Q})=\emptyset$.
Proof.
(1) follows from the definition and Lemma 5.3. Since $\langle P,(N/M)P\rangle=f_{N}(t(E,P))\otimes(1/M)$ by Proposition 4.2, we get (2) from (1) together with an elementary observation that for $z\in\mathbb{Q}^{\times}$ one has $z\otimes(1/M)=0$ in $\mathbb{Q}^{\times}\otimes(\mathbb{Q}/\mathbb{Z})$ if and only if $z=\pm s^{M}$ for some $s\in\mathbb{Q}^{\times}$. To see (3), it suffices to note that there is a finite morphism $X_{1}(N)_{M}^{\epsilon}\to X_{1}(N)_{M^{\prime}}^{\epsilon}$ induced by $s\mapsto s^{M/M^{\prime}}$. We are done. ∎
Corollary 5.5.
For $(E,P)\in Y_{1}(N)(\mathbb{Q})$, we have $(E(\mathbb{Q})_{\operatorname{tors}},E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}})=(\langle P\rangle,\langle(N/M)P\rangle)$ if and only if all of the following conditions are met:
  1. (1)
    $(E,P)\not\in\operatorname{Im}(Y_{1}(N\ell)(\mathbb{Q})\to Y_{1}(N)(\mathbb{Q}))$ for any $\ell\in\{2,3,5\}$ such that $N\ell\leq 12$.
  2. (2)
    $(E,P)\not\in\operatorname{Im}(Y_{1}^{0}(N,2)(\mathbb{Q})\to Y_{1}(N)(\mathbb{Q}))$.
  3. (3)
    $(E,P)\in\operatorname{Im}(Y_{1}(N)_{M}^{\pm}(\mathbb{Q})\to Y_{1}(N)(\mathbb{Q}))$.
  4. (4)
    $(E,P)\not\in\operatorname{Im}(Y_{1}(N)_{M^{\prime}}^{\pm}(\mathbb{Q})\to Y_{1}(N)(\mathbb{Q}))$ for any $M^{\prime}>M$ such that $M|M^{\prime}|N$.
Here $Y_{1}^{0}(N,2)$ denotes the complement of cusps in $X_{1}^{0}(N,2)$.
Proof.
(1) and (2) mean $E(\mathbb{Q})_{\operatorname{tors}}=\langle P\rangle$. Under these conditions, (3) and (4) mean $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=\langle(N/M)P\rangle$ by Lemma 5.4 (2). ∎
These modular curves turns out to be models of the classical modular curves.
Theorem 5.6.
Assume that $N\in\{4,5,6,7,8,9,10,12\}$ and let $M$ be a positive divisor of $N$. Denote by $X_{1}(MN,N)$ the modular curve over $\mathbb{Q}$ associated to the congruence subgroup
$\Gamma_{1}(MN,N):=\Gamma_{0}(MN)\cap\Gamma_{1}(N)=\left\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in{\operatorname{SL}}(2,\mathbb{Z}):\ a,d\equiv 1\bmod N,\ MN|c\right\},$
on which the cusp 0 is $\mathbb{Q}$-rational. Then $X_{1}(N)_{M}^{+}$ is isomorphic to $X_{1}(MN,N)$ over $\mathbb{Q}$.
Proof.
Let
$s(\tau)=\left(\frac{E_{0,1}^{(N)}(\tau)E_{0,3}^{(N)}(\tau)}{E_{0,2}^{(N)}(\tau)^{2}}\right)^{N/M}.$
By Lemma 3.3, for $\gamma=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)\in\Gamma_{1}(N)$, we have
$\begin{split}s(\gamma\tau)&=\left(e^{2\pi icd/N^{2}}\frac{E_{c,d}^{(N)}(\tau)E_{3c,3d}^{(N)}(\tau)}{E_{2c,2d}^{(N)}(\tau)^{2}}\right)^{N/M}=\left(e^{2\pi icd/N^{2}}\frac{E_{c,1}^{(N)}(\tau)E_{3c,3}^{(N)}(\tau)}{E_{2c,2}^{(N)}(\tau)^{2}}\right)^{N/M}\\ &=\left(e^{2\pi icd/N^{2}}e^{-4\pi ic/N^{2}}\frac{E_{0,1}^{(N)}(\tau)E_{0,3}^{(N)}(\tau)}{E_{0,2}^{(N)}(\tau)^{2}}\right)^{N/M}=e^{-2\pi ic/MN}s(\tau).\end{split}$
Therefore, $s(\gamma\tau)=s(\tau)$ if $\gamma\in\Gamma_{1}(MN,N)$. Also, $\mathbb{C}(s(\tau),t(\tau)):\mathbb{C}(t(\tau))]=M=[\mathbb{C}(X_{1}(MN,N)):\mathbb{C}(X_{1}(N))]$. Consequently, $s(\tau)$ and $t(\tau)$ generate the field of modular functions on $X_{1}(MN,N)$. Moreover, using the transformation formula for generalized Dedekind eta functions, we see that the coefficients in the Fourier expansion of $s(\tau)$ at the cusp 0 are all rational numbers. Thus, the relation $s^{M}=f_{N}(t)$ is an equation for $X_{1}(MN,N)$ over $\mathbb{Q}$. This completes the proof of the theorem. ∎
Corollary 5.7.
If $Y_{0}(MN)(\mathbb{Q})=\emptyset$, then $Y_{1}(N)_{M}^{+}(\mathbb{Q})=\emptyset$. (According to Mazur and Kenku [11, 6] (see also [2, Theorem 2.2]), one has $Y_{0}(N)(\mathbb{Q})=\emptyset$ precisely when $N\geq 20$ and $N\not\in\{21,25,27,37,43,67,163\}$.)
Proof.
This is a direct consequence of Theorem 5.6, since there is a morphism $X_{1}(MN,N)\to X_{0}(MN)$ over $\mathbb{Q}$ (as $\Gamma_{1}(MN,N)\subset\Gamma_{0}(MN)$ by definition). ∎
Remark 5.8.
Let $k$ be a subfield of $\mathbb{C}$. Then $X_{1}(MN,N)(k)$ parameterize isomorphism classes of triples $(E,P,C)$ of an elliptic curve $E$ over $k$, a $k$-rational point $P$ of order $N$, and a cyclic subgroup $C$ of order $MN$ defined over $k$, subject to the condition $MC=\langle P\rangle$. Given such a triple $(E,P,C)$, take $\tau\in\mathbb{H}$ such that $(E,P,C)\cong(\mathbb{C}/(\mathbb{Z}\tau+\mathbb{Z}),1/N,\langle 1/MN\rangle)$, and put
$s:=\left(\frac{E_{0,1}^{(N)}(\tau)E_{0,3}^{(N)}(\tau)}{E_{0,2}^{(N)}(\tau)^{2}}\right)^{N/M}.$
Then, $(E,P,C)$ corresponds to $(E,P,s)\in X_{1}^{0}(N)_{M}^{+}(k)$ under the the isomorphism in Theorem 5.6. This can be read off from the proof of Theorem 5.6.
Proposition 5.9.
Assume that $N\in\{4,5,6,7,8,9,10,12\}$ and let $M$ be a positive divisor of $N$, and $\epsilon\in\{+,-\}$.
  1. (1)
    If one of the following conditions holds, $X_{1}(N)_{M}^{\epsilon}$ is isomorphic to $\mathbb{P}^{1}$ over $\mathbb{Q}$.
    1. (a)
      $M=1$.
    2. (b)
      $(N,M)\in\{(4,2),(4,4),(5,5),(6,2),(6,3),(8,2)\}$.
  2. (2)
    If neither of (a), (b) is satisfied, then we have $Y_{1}(N)_{M}^{\epsilon}(\mathbb{Q})=\emptyset$.
Proof.
(1) Since $X_{1}(N)_{M}^{\epsilon}$ always has a $\mathbb{Q}$-rational point by Remark 5.2 (2), it suffices to show that its genus is zero. For this, we may assume $\epsilon=+$ by Remark 5.2 (1), but then Theorem 5.6 shows $X_{1}(N)_{M}^{+}$ is isomorphic to $X_{1}(MN,N)$, which has genus zero. (This can also be seen easily from the defining equation $s^{M}=f_{N}(z)$.)
(2) We first note that the cases $(N,M)\in\{(8,8),(9,9),(10,10),(12,4),(12,6),(12,12)\}$ are reduced to the case $(N,M^{\prime})$ for some smaller $M^{\prime}|N$ by Lemma 5.4 (3). Next, if $\epsilon=+$ or $M$ is odd, we can apply Remark 5.2 (1) and Corollary 5.7 to deduce $Y_{1}(N)_{M}^{\epsilon}(\mathbb{Q})=\emptyset$. This covers the cases $(M,N)\in\{(7,7),(10,5),(12,3)\}$ with $\epsilon\in\{+,-\}$, and $(M,N)\in\{(6,6),(8,4),(10,2),(12,2)\}$ with $\epsilon=$-.
This leaves us with the cases $(N,M)\in\{(6,6),(8,4),(9,3),(10,2),(12,2)\}$ with $\epsilon=$-. (For $(N,M)=(9,3)$, the value of $\epsilon$ does not matter because of Remark 5.2 (1).) In all cases, there exists an isomorphism $\alpha:E\to X_{1}(N)_{M}^{\epsilon}$, where $E$ is an elliptic curve over $\mathbb{Q}$ with finitely many $\mathbb{Q}$-rational points, as exhibited in the following table, where we set $h_{1}(x)=(1-x)(1+x)^{4}(1+x-x^{2})^{2}$ and $h_{2}(x)=(1+x)(1+x+x^{2})(1+x^{2})^{2}(1-x)^{3}$.
[…]
Since all points of $\alpha(E(\mathbb{Q}))$ are cusps by Lemma 3.6, we get $Y_{1}(\mathbb{Q})^{\epsilon}_{M}=\emptyset$ in these cases as well. We are done. ∎
5.2. Modular curve $X_{1}^{0}(N,2)_{\underline{M}}^{\pm}$
Definition 5.10.
Assume $N\in\{4,6,8\}$. We set
$\mathscr{M}_{N}:=\{\underline{M}=(M_{1},M_{2},M_{3})\mid M_{1}|N,\ M_{2},M_{3}\in\{1,2\}\},\$
$\mathscr{S}:=\{+,-\}^{3}=\{\underline{\epsilon}=(\epsilon_{1},\epsilon_{2},\epsilon_{3})\mid\epsilon_{1},\epsilon_{2},\epsilon_{3}\in\{+,-\}\}.$
Let $g_{N}^{(\nu)}(u)\ (\nu=1,2,3)$ be the function from Proposition 4.3. For $\underline{M}\in\mathscr{M}_{N}$ and $\underline{\epsilon}\in\mathscr{S}$, we define $X_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}$ to be the smooth projective model over $\mathbb{Q}$ of the curve defined by $\epsilon_{i}v_{\nu}^{M_{\nu}}=g_{N}^{(\nu)}(u)\ (\nu=1,2,3)$. In other words, $X_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}$ is the normalization of the fiber product
${X_{1}^{0}}^{\prime}(N,2)_{\underline{M}}^{\underline{\epsilon}}:=X_{1}^{0}(N,2)\times_{(\mathbb{P}^{1}_{w_{1}}\times\mathbb{P}^{1}_{w_{2}}\times\mathbb{P}^{1}_{w_{3}})}(\mathbb{P}^{1}_{v_{1}}\times\mathbb{P}^{1}_{v_{2}}\times\mathbb{P}^{1}_{v_{3}}),$
where $X_{1}^{0}(N,2)\to\mathbb{P}^{1}_{w_{1}}\times\mathbb{P}^{1}_{w_{2}}\times\mathbb{P}^{1}_{w_{3}}$ is the composition of the isomorphism $X_{1}^{0}(N,2)\cong\mathbb{P}^{1}_{u}$ from Lemma 3.7 followed by $\mathbb{P}^{1}_{u}\to\mathbb{P}^{1}_{w_{1}}\times\mathbb{P}^{1}_{w_{2}}\times\mathbb{P}^{1}_{w_{3}},\ w_{\nu}=g_{N}^{(\nu)}(u)\ (\nu=1,2,3)$, and $\mathbb{P}^{1}_{v_{1}}\times\mathbb{P}^{1}_{v_{2}}\times\mathbb{P}^{1}_{v_{3}}\to\mathbb{P}^{1}_{w_{1}}\times\mathbb{P}^{1}_{w_{2}}\times\mathbb{P}^{1}_{w_{3}}$ is defined by $w_{\nu}=v_{\nu}^{M_{\nu}}\ (\nu=1,2,3)$. (Here $\times$ denotes the fiber product over $\operatorname{Spec}\mathbb{Q}$.) We write $Y_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}\subset X_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}$ for the inverse image of the complement $Y_{1}^{0}(N,2)$ of cusps in $X_{1}^{0}(N,2)$. Also, we write $X_{1}^{0}(N,2)_{\underline{M}}^{\pm}$ for the disjoint union of $X_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}$ over $\underline{\epsilon}\in\mathscr{S}$, and similarly for $Y_{1}^{0}(N,2)_{\underline{M}}^{\pm}$.
Remark 5.11.
  1. (1)
    Observe that $u=0$ is a simple zero of $g_{N}^{(1)}(u)$, and $g_{N}^{(\nu)}(0)=1\in\mathbb{P}^{1}_{w_{\nu}}$ is not a branch point of $\mathbb{P}^{1}_{v_{\nu}}\to\mathbb{P}^{1}_{w_{\nu}},w_{\nu}=v_{\nu}^{M_{\nu}}$ for $\nu=2,3$. It follows that, if $\epsilon_{2}=\epsilon_{3}=+$, then ${X_{1}^{0}}^{\prime}(N)_{\underline{M}}^{\underline{\epsilon}}$ is non-singular at $(u,v_{1},v_{2},v_{3})=(0,0,1,1)\in{X_{1}^{0}}^{\prime}(N,2)_{\underline{M}}^{\underline{\epsilon}}(\mathbb{Q})$, and hence there is a unique $\mathbb{Q}$-rational point on $X_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}$ above it.
  2. (2)
    For $\underline{M}=(M_{1},M_{2},M_{3}),\underline{M}^{\prime}=(M_{1}^{\prime},M_{2}^{\prime},M_{3}^{\prime})\in\mathscr{M}_{N}$, we write $\underline{M}^{\prime}|\underline{M}$ if $M_{\nu}^{\prime}|M_{\nu}$ holds for all $\nu=1,2,3$. In this case, there is a finite morphism
    $X_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}\to X_{1}^{0}(N,2)_{\underline{M}^{\prime}}^{\underline{\epsilon}}$
    induced by $v_{\nu}\mapsto v_{\nu}^{M_{\nu}/M_{\nu}^{\prime}}\ (\nu=1,2,3)$. By taking $\underline{M}^{\prime}=(1,1,1)$ we get
    $X_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}\to X_{1}^{0}(N,2)_{(1,1,1)}^{\underline{\epsilon}}=X_{1}^{0}(N,2).$
Lemma 5.12.
The canonical map $Y_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}\to{X_{1}^{0}}^{\prime}(N,2)_{\underline{M}}^{\underline{\epsilon}}$ is an isomorphism onto its image.
Proof.
This is shown by the same argument as Lemma 5.3. ∎
Recall that we have specified an isomorphism $u:X_{1}^{0}(N,2)\cong\mathbb{P}^{1}$ in Lemma 3.7. By the moduli interpretation, there is a bijection between $Y_{1}^{0}(N,2)(\mathbb{Q})$ and the set of isomorphism classes of triples $(E,P,Q)$ of an elliptic curve $E$ over $\mathbb{Q}$, a $\mathbb{Q}$-rational point $P$ of order $N$, and a $\mathbb{Q}$-rational point $Q$ of order $2$ such that $Q\not\in\langle P\rangle$. By abuse of notation, we write $(E,P,Q)\in Y_{1}^{0}(N,2)(\mathbb{Q})$ in this situation, and write $u(E,P,Q)\in\mathbb{Q}^{\times}\subset\mathbb{P}^{1}(\mathbb{Q})$ for the point corresponding to $(E,P,Q)\in Y_{1}^{0}(N,2)(\mathbb{Q})$.
Lemma 5.13.
Let $N,\underline{M},\underline{\epsilon}$ be as in Definition 5.10.
  1. (1)
    There is a bijection between $Y_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}(\mathbb{Q})$ and the set of sextuple $(E,P,Q,v_{1},v_{2},v_{3})$ where $(E,P,Q)\in Y_{1}^{0}(N,2)(\mathbb{Q})$ and $v_{\nu}\in\mathbb{Q}^{\times}$ such that $\epsilon_{\nu}v_{\nu}^{M_{\nu}}=g_{N}^{(\nu)}(u(E,P,Q))$ $(\nu=1,2,3)$.
  2. (2)
    Given $(E,P,Q)\in Y_{1}^{0}(N,2)(\mathbb{Q})$, we have
    $\langle P,\frac{N}{M_{1}}P\rangle=\langle Q,\frac{2}{M_{2}}Q\rangle=\langle P,\frac{2}{M_{3}}Q\rangle=0$
    if and only if $(E,P,Q)$ belongs to the image of the canonical map $Y_{1}^{0}(N,2)_{\underline{M}}^{\pm}(\mathbb{Q})\to Y_{1}^{0}(N,2)(\mathbb{Q})$.
  3. (3)
    Suppose $\underline{M}^{\prime}|\underline{M}$. If $X_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}\cong\mathbb{P}^{1}$, then $X_{1}^{0}(N,2)_{\underline{M}^{\prime}}^{\underline{\epsilon}}\cong\mathbb{P}^{1}$. If $Y_{1}^{0}(N,2)_{\underline{M}^{\prime}}^{\underline{\epsilon}}(\mathbb{Q})=\emptyset$, then $Y_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}(\mathbb{Q})=\emptyset$.
Proof.
Again, this is shown by the same argument as Lemma 5.4. ∎
Corollary 5.14.
Suppose $(E,P,Q)\in Y_{1}^{0}(N,2)(\mathbb{Q})$ and $E(\mathbb{Q})_{\operatorname{tors}}=\langle P,Q\rangle$. Let $M$ be a positive divisor of $N$, and put $m:=\gcd(N/M,2)$. Then we have
$\frac{N}{M}P\in E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}\Leftrightarrow(E,P,Q)\in\operatorname{Im}(Y_{1}^{0}(N,2)_{(M,1,m)}^{\pm}(\mathbb{Q})\to Y_{1}^{0}(N,2)(\mathbb{Q})),$
$Q\in E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}\Leftrightarrow(E,P,Q)\in\operatorname{Im}(Y_{1}^{0}(N,2)_{(1,2,2)}^{\pm}(\mathbb{Q})\to Y_{1}^{0}(N,2)(\mathbb{Q})).$
Proof.
This follows immediately from Lemma 5.13. ∎
Here is an analogue of Proposition 5.9.
Proposition 5.15.
Let $N,\underline{M},\underline{\epsilon}$ be as in Definition 5.10.
  1. (1)
    If one of the following conditions holds, $X_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}$ is isomorphic to $\mathbb{P}^{1}$ over $\mathbb{Q}$.
    1. (a)
      $N=4$ and $\underline{M}\in\{(2,1,2),(1,2,2),(2,1,1),(1,2,1),(1,1,2),(1,1,1)\}$.
    2. (b)
      $N=6$ and $\underline{M}\in\{(1,1,2),(1,1,1)\}$.
    3. (c)
      $N=8$ and $\underline{M}=(1,1,1)$.
  2. (2)
    If none of (a)–(c) is satisfied, then we have $Y_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}(\mathbb{Q})=\emptyset$.
Proof.
(1a) It suffices to show the first two cases by Lemma 5.13 (3). Since $g_{4}^{(3)}(u)=1+u$ defines an isomorphism $\mathbb{P}^{1}_{u}\to\mathbb{P}^{1}_{w_{3}}$, we have $X_{1}^{0}(4,2)_{(2,1,2)}^{\underline{\epsilon}}\cong X_{1}^{0}(4,2)_{(2,1,1)}^{\underline{\epsilon}}$ and $X_{1}^{0}(4,2)_{(1,2,2)}^{\underline{\epsilon}}\cong X_{1}^{0}(4,2)_{(1,2,1)}^{\underline{\epsilon}}$. Now ${X_{1}^{0}}^{\prime}(N,2)_{(2,1,1)}^{\underline{\epsilon}}$ is a nodal curve $4u(1+u)^{2}=\pm v_{1}^{2}$, and ${X_{1}^{0}}^{\prime}(N,2)_{(1,2,1)}^{\underline{\epsilon}}$ is a conic $1-u^{2}=\pm v_{2}^{2}$ having a $\mathbb{Q}$-rational point, both of which become isomorphic to $\mathbb{P}^{1}$ after normalization.
(1b) It suffices to show the first case by Lemma 5.13 (3). Now $X_{1}^{0}(6,2)_{(1,1,2)}^{\underline{\epsilon}}$ is a conic defined by $(1-u)(1-3u)=\pm v_{3}^{2}$ having a $\mathbb{Q}$-rational point, which is rational.
(1c) This is obvious since ${X_{1}^{0}}(8,2)_{(1,1,1)}^{\underline{\epsilon}}\cong{X_{1}^{0}}(8,2)$.
(2) We divide the proof by the value of $N$.
Case of $N=4$
It suffices to show the cases $\underline{M}=(4,1,1),(2,2,1)$ by Lemma 5.13 (3).
  • Suppose first $\underline{M}=(4,1,1)$ so that ${X_{1}^{0}}(4,2)_{\underline{M}}^{\underline{\epsilon}}$ is given by $4u(1+u)^{2}=\pm v_{1}^{4}$. It is isomorphic to an elliptic curve $E_{\pm}:y^{2}=x^{3}\pm 4x$ by the change of variables $(u,v_{1})=(-1\pm(y^{2}/4x),\pm y/2)$ ($(x,y)=(0,0)$ corresponds to $(u,v_{1})=(-1,0)$). As $E_{+}(\mathbb{Q})=\{(0,0),(2,\pm 4),O\}$ and $E_{-}(\mathbb{Q})=\{(0,0),(\pm 2,0),O\}$, we find all points of ${X_{1}^{0}}(4,2)_{\underline{M}}^{\underline{\epsilon}}(\mathbb{Q})$ are cusps.
  • Next suppose $\underline{M}=(2,2,1)$ so that ${X_{1}^{0}}(4,2)_{\underline{M}}^{\underline{\epsilon}}$ is given by $4u(1+u)^{2}=\pm v_{1}^{2},\ 1-u^{2}=\pm v_{2}^{2}$. There is a finite morphism from ${X_{1}^{0}}(4,2)_{\underline{M}}^{\underline{\epsilon}}$ onto an elliptic curve $E:y^{2}=x^{3}\pm x$ given by $(x,y)=(\mp u,(v_{1}v_{2})/2)$. As $E(\mathbb{Q})=\{(0,0),(\pm 1,0),O\}$, we conclude all rational points on ${X_{1}^{0}}(4,2)_{\underline{M}}^{\underline{\epsilon}}$ are cusps.
Case of $N=6$
It suffices to consider the cases $\underline{M}=(3,1,1),(2,1,1),(1,2,1)$.
  • Suppose first $\underline{M}=(3,1,1)$. We may assume $\epsilon_{1}=+$ (as $M_{1}=3$ is odd), and then ${X_{1}^{0}}(6,2)_{\underline{M}}^{\underline{\epsilon}}$ is given by $u(1-u)(1+3u)^{3}(1+u)^{4}=v_{1}^{3}$. It is isomorphic to an elliptic curve $E:x-x^{3}=y^{3}$ by the change of variables $(x,y)=(u,v_{1}/((1+3u)(1+u)))$. All points of $E(\mathbb{Q})=\{(0,1),(\pm 1,0),(\pm 1/3,\pm 2/3),O\}$ correspond to cusps.
  • Suppose $\underline{M}=(2,1,1)$ so that ${X_{1}^{0}}(6,2)_{\underline{M}}^{\underline{\epsilon}}$ is given by $u(1-u)(1+3u)^{3}(1+u)^{4}=\pm v_{1}^{2}$. It is isomorphic to an elliptic curve $E_{\pm}:x(1-x)(1+3x)=\pm y^{2}$ by $(x,y)=(u,v_{1}/((1+3u)(1+u)^{2}))$. Again, all points of
    $E_{+}(\mathbb{Q})=\{(0,0),(1,0),(-1,\pm 2),(-1/3,0),(1/3,\pm 2/3),O\},$
    $E_{-}(\mathbb{Q})=\{(0,0),(1,0),(-1/3,0),O\}$
    correspond to cusps.
  • Finally, suppose $\underline{M}=(1,2,1)$ so that ${X_{1}^{0}}(6,2)_{\underline{M}}^{\underline{\epsilon}}$ is given by $(1-u^{2})(1-9u^{2})=\pm v_{2}^{2}$. We find ${X_{1}^{0}}(6,2)_{\underline{M}}^{\underline{\epsilon}}(\mathbb{Q})=\{(\pm 1,0),(\pm 1/3,0),(0,\pm 1),\pm\infty\}$ if $\epsilon_{2}=+$, and ${X_{1}^{0}}(6,2)_{\underline{M}}^{\underline{\epsilon}}(\mathbb{Q})=\{(\pm 1,0),(\pm 1/3,0)\}$ if $\epsilon_{2}=$- (in $(u,v_{2})$-coordinates). Again, these points correspond to cusps.
Case of $N=8$
It suffices to consider the cases $\underline{M}=(2,1,1),(1,2,1),(1,1,2)$.
  • Suppose first $\underline{M}=(2,1,1)$. By Lemma 5.13 (2), $(E,P,Q)\in Y_{1}^{0}(8,2)(\mathbb{Q})$ is in the image from ${Y_{1}^{0}}(8,2)_{\underline{M}}^{\pm}(\mathbb{Q})$ if and only if $\langle P,4P\rangle=0$. This implies $(E,2P,Q)\in Y_{1}^{0}(4,2)(\mathbb{Q})$ is in the image from ${Y_{1}^{0}}(4,2)_{(4,1,1)}^{\pm}(\mathbb{Q})$, but we have already shown the latter set is empty.
  • Next suppose $\underline{M}=(1,2,1)$ so that ${X_{1}^{0}}(8,2)_{\underline{M}}^{\underline{\epsilon}}$ is a genus one curve defined by $1-6u^{2}+u^{4}=\pm v_{2}^{2}$. We have ${X_{1}^{0}}(8,2)_{\underline{M}}^{\underline{\epsilon}}=\{(0,\pm 1),\pm\infty\}$ if $\epsilon_{2}=+$, and ${X_{1}^{0}}(8,2)_{\underline{M}}^{\underline{\epsilon}}=\{(\pm 1,\pm 2),\pm\infty\}$ if $\epsilon_{2}=$- (in $(u,v_{2})$-coordinates), all of which are cuspidal.
  • Finally, suppose $\underline{M}=(1,1,2)$ so that ${X_{1}^{0}}(8,2)_{\underline{M}}^{\underline{\epsilon}}$ is a hyperelliptic curve defined by $(1-u^{4})(1+2u-u^{2})=\pm v_{3}^{2}$. We find ${X_{1}^{0}}(8,2)_{\underline{M}}^{\underline{\epsilon}}(\mathbb{Q})=\{(\pm 1,0),(0,\pm 1),\pm\infty\}$ if $\epsilon_{3}=+$, and ${X_{1}^{0}}(8,2)_{\underline{M}}^{\underline{\epsilon}}(\mathbb{Q})=\{(\pm 1,0)\}$ if $\epsilon_{3}=$- (in $(u,v_{3})$-coordinates). Again, these points are cuspidal. ∎
6. End of the proof
We are now in position to complete the proof of Theorem 1.1.
6.1. Proof of non-existence
We first consider the case $E(\mathbb{Q})_{\operatorname{tors}}$ is cyclic. Let $E$ be an elliptic curve over $\mathbb{Q}$, $A=\mathbb{Z}/N\mathbb{Z}$ with $N\in\{6,7,8,9,10,12\}$, and $\alpha:E(\mathbb{Q})_{\operatorname{tors}}\cong A$ an isomorphism. Take a generator $P\in E(\mathbb{Q})_{\operatorname{tors}}$ and a positive divisor $M|N$ such that $M$ does not appear as the order of $B$ in the list of Theorem 1.1. If $(N/M)P\in E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}$ would hold, Lemma 5.4 implies $(E,P)\in\operatorname{Im}(Y_{1}(N)_{M}^{\pm}(\mathbb{Q})\to Y_{1}(N)(\mathbb{Q}))$, but we have $Y_{1}(N)_{M}^{\pm}(\mathbb{Q})=\emptyset$ by Proposition 5.9. This proves that there is no $E$ such that $(E(\mathbb{Q})_{\operatorname{tors}},E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}})=(\langle P\rangle,\langle(N/M)P\rangle)$.
The proofs of the cases $A=\mathbb{Z}/N\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ for $N=4,6,8$ are similar. As a sample, we detail the case $N=6$. Let $E$ an elliptic curve over $\mathbb{Q}$, and let $\alpha:E(\mathbb{Q})_{\operatorname{tors}}\cong A:=\mathbb{Z}/6\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ be an isomorphism. If $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}$ were non-trivial, there should exist $P,Q\in E(\mathbb{Q})_{\operatorname{tors}}$ such that $(E,P,Q)\in Y_{1}^{0}(6,2)(\mathbb{Q})$ and $\{2P,3P,Q\}\cap E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}\not$$\emptyset$. Since we have $Y_{1}^{0}(6,2)_{(3,1,1)}^{\pm}(\mathbb{Q})=Y_{1}^{0}(6,2)_{(2,1,2)}^{\pm}(\mathbb{Q})=Y_{1}^{0}(6,2)_{(1,2,2)}^{\pm}(\mathbb{Q})=\emptyset$ by Proposition 5.15, this would contradict Lemma 5.13 (2). The cases $N=4,6$ are similar and skipped.
The reason for the non-existence in the case $A=\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ is quite different. Indeed, there do exist infinitely many $(E,P,Q)\in Y_{1}^{0}(2,2)(\mathbb{Q})$ with $\langle P,P\rangle=\langle Q,Q\rangle=\langle P,Q\rangle=0$. However, all of them come from $Y_{1}^{0}(4,2)(\mathbb{Q})$, as is shown in the following proposition, which completes the proof of the non-existence part of Theorem 1.1.
Proposition 6.1.
Let $(E,P,Q)\in Y_{1}^{0}(2,2)(\mathbb{Q})$ and suppose that $P,Q\in E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}$. Then $E(\mathbb{Q})_{\operatorname{tors}}$ contains an element of order four. Consequently, there is no elliptic curve $E$ over $\mathbb{Q}$ such that $E(\mathbb{Q})_{\operatorname{tors}}=E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}\simeq\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$.
Proof.
We may assume $E$ is given by $E_{2,a,u}^{\prime}:y^{2}=x(x-a)(x-au)$ from (4.4) for some $a,u\in\mathbb{Q}$ with $au(1-u)\not=0$, and $P=(0,0),Q=(a,0)$. By Lemma 4.7, we have $\langle P,P\rangle=u\otimes(1/2)$, $\langle Q,Q\rangle=(1-u)\otimes(1/2)$ and $\langle P,Q\rangle=a\otimes(1/2)$. The assumption $P,Q\in E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}$ implies that $u=\pm b^{2}$, $1-u=\pm c^{2}$, and $a=\pm d^{2}$ for some $b,c,d\in(\mathbb{Q}^{\times})^{2}$. Recall from Example 6.2 of [5] that an elliptic curve $y^{2}=x(x-A)(x-B)\ (A,B\in\mathbb{Q})$ has a rational torsion point $R$ of order $4$ such that $2R=(0,0)$ if and only if $A$ and $B$ are squares in $\mathbb{Q}^{\times}$. In the case $(u,1-u,a)=(b^{2},\pm c^{2},d^{2})$, the elliptic curve $E_{2,a,u}^{\prime}$ is $y^{2}=x(x-d^{2})(x-b^{2}d^{2})$, and hence it has a torsion point of order $4$. In the case $(u,1-u,a)=(-b^{2},c^{2},d^{2})$, we make a change a variable $(x,y)\mapsto(x-b^{2}d^{2},y)$ in $E_{2,a,u}^{\prime}$ and find that $E_{2,a,u}^{\prime}$ is isomorphic to $y^{2}=x(x-b^{2}d^{2})(x-c^{2}d^{2})$. Again, this elliptic curve has a rational point of order $4$. When $(u,1-u,a)=(-b^{2},c^{2},-d^{2})$, we make a change of variables $(x,y)\mapsto(x-d^{2},y)$ and find that $E_{2,a,u}^{\prime}$ is isomorphic to $y^{2}=x(x-d^{2})(x-c^{2}d^{2})$, which has a rational point of order $4$ as well. ∎
6.2. Proof of existence
First let us consider the cases $A=\mathbb{Z}/N\mathbb{Z}$ with $N=1,7,9,10,12$ or $A=\mathbb{Z}/N\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ with $N=6,8$. We have shown that $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}$ must be trivial if $E(\mathbb{Q})_{\operatorname{tors}}\cong A$. Since there are infinitely many $E$ with $E(\mathbb{Q})_{\operatorname{tors}}\cong A$ by Mazur’s theorem, the existence part of Theorem 1.1 follows in these cases.
To proceed, we use the following fact about the thin sets (see [14, Proposition 3.4.2]).
Proposition 6.2.
If $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ is a finite morphism of degree $d$, then the number of points of $f(\mathbb{P}^{1}(\mathbb{Q}))\subset\mathbb{P}^{1}(\mathbb{Q})$ with height $\leq h$ is asymptotically $\sim h^{2/r}$ as $h\to\infty$.
For convenience, we describe the degrees of relevant morphisms.
${X_{1}(12)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{8}$$\scriptstyle{4}$${X_{1}(4)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{2}$${X_{1}(8)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{4}$${X_{1}(6)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{4}$$\scriptstyle{3}$${X_{1}(2)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{3}$${X_{1}(10)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{12}$$\scriptstyle{3}$${X_{1}(9)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{9}$${X_{1}(3)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{4}$${X_{1}(1)}$${X_{1}(5)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{12}$${X_{1}^{0}(6,2)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{4}$$\scriptstyle{2}$${X_{1}^{0}(2,2)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{2}$${X_{1}^{0}(4,2)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{2}$$\scriptstyle{2}$${X_{1}^{0}(8,2)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{2}$$\scriptstyle{4}$${X_{1}(6)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{4}$${X_{1}(2)}$${X_{1}(4)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{2}$${X_{1}(8)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{4}$
Recall also that the degree of $X_{1}(N)_{M}^{\epsilon}\to X_{1}(N)$ is $M$, and that of $X_{1}^{0}(N,2)_{\underline{M}}^{\underline{\epsilon}}\to X_{1}^{0}(N,2)$ is $M_{1}M_{2}M_{3}$. Our proof goes from a bigger groups to smaller.
Case of $A=\mathbb{Z}/4\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$
Let $(E,P,Q)\in Y_{1}^{0}(4,2)(\mathbb{Q})$, and suppose first $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}$ is non-trivial. Then we must have $E(\mathbb{Q})_{\operatorname{tors}}\cong A$ (because we have shown $E(\mathbb{Q})_{\operatorname{tors}}\cong\mathbb{Z}/8\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ implies $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=0$), and then $|E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}|=2$ (because we have shown $|E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}|\not=4,8$). Thus we have
  • $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=\langle 2P\rangle\Leftrightarrow(E,P,Q)\in\operatorname{Im}(Y_{1}^{0}(4,2)_{(2,1,1)}^{\pm}(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q}))$,
  • $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=\langle Q\rangle\Leftrightarrow(E,P,Q)\in\operatorname{Im}(Y_{1}^{0}(4,2)_{(1,2,2)}^{\pm}(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q}))$.
Since the components of $Y_{1}^{0}(N)_{(2,1,1)}^{\pm}$ and $Y_{1}^{0}(N)_{(1,2,2)}^{\pm}$ are isomorphic to $\mathbb{P}^{1}$ by Proposition 5.15, we conclude by Proposition 6.2 that there are infinitely many such $(E,P,Q)$ in both cases. Finally, we have $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=0$ precisely when neither of the above conditions are met and $(E,P,Q)\not\in\operatorname{Im}(Y_{1}^{0}(8,2)(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q}))$. Again we conclude by Proposition 6.2 that there are infinitely many such $(E,P,Q)$.
Case of $A=\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$
We need a special treatment for this case. Consider the following subsets of $Y_{1}(1)(\mathbb{Q})$.
$J:=$ $\operatorname{Im}(Y_{1}(2,2)(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
$J_{1}:=$ $\operatorname{Im}(Y_{1}(4,2)(\mathbb{Q})\sqcup Y_{1}(6,2)(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
$J_{2}:=$ $\operatorname{Im}(Y_{1}(4,2)_{(2,1,1)}^{\pm}(\mathbb{Q})\sqcup Y_{1}(4,2)_{(1,2,2)}^{\pm}(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
$J_{3}:=$ $\{j(E)\in Y_{1}(1)(\mathbb{Q})\mid(E,P,Q)\in Y_{1}(2,2)(\mathbb{Q}),\ E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}\not=0\}.$
By Corollary 5.14 and the (proved) non-existence result, we obtain
  • $J\setminus(J_{1}\cup J_{3})$ is precisely the set of the $j$-invariants of elliptic curves $E$ over $\mathbb{Q}$ such that $E(\mathbb{Q})_{\operatorname{tors}}\cong A$ and $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=0$.
  • $J_{3}\setminus J_{2}$ is precisely the set of the $j$-invariants of elliptic curves $E$ over $\mathbb{Q}$ such that $E(\mathbb{Q})_{\operatorname{tors}}\cong A$ and $|E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}|=2$.
Furthermore, we claim that there is a morphism $\widetilde{j}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ of degree six such that $J_{3}=\widetilde{j}(\mathbb{P}^{1}(\mathbb{Q})\setminus\{0,\pm 1,\infty\})$. By Proposition 6.2, this will complete the proof of this case.
To show the claim, we make use of the elliptic curve $E_{2,a,u}^{\prime}:y^{2}=x(x-a)(x-au)$ from (4.4) for $a,u\in\mathbb{Q}$ with $au(1-u)\not=0$ again. Observe that its $j$-invariant is given by
$j(u)=256\frac{(1-u+u^{2})^{3}}{u^{2}(1-u)^{2}}.$
It suffices to show $J_{3}=\{j(\pm v^{2})\mid v\in\mathbb{Q},v\not=0,\pm 1\}$. Let $(E,P,Q)\in Y_{1}^{0}(2,2)(\mathbb{Q})$ and suppose $P\in E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}$. There is an isomorphism $E\cong E_{2,a,u}^{\prime}$ for some $a,u\in\mathbb{Q}$ with $au(1-u)\not=0$, by which $P$ corresponds to $(0,0)$. Since $P\in E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}$, we have $u=\pm v^{2}$ for some $v\in\mathbb{Q}\setminus\{0,\pm 1\}$ by Lemma 4.7. Hence the $j(u)=j(\pm v^{2})$ belongs to $J_{3}$. The converse is seen by taking $u=\pm v^{2}$ (and $a=1$). This completes the proof of the claim and hence the existence part of this case.
Cyclic case
In view of Proposition 6.2, the cases $A=\mathbb{Z}/N\mathbb{Z},N=4,5,6,8$ now follow from Corollary 5.5. The remaining cases $N=2,3$ are proved by the same way as $A=\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$, using the formulas
$j(E_{2,t,a})=\begin{cases}1024\frac{(t+1)^{2}(t+4)}{a^{4}t^{2}},&(t\not=-1),\\ 1728&(t=-1),\end{cases}\quad\begin{cases}j(E_{3,t})=27\frac{(t+1)(t+9)^{3}}{t^{3}}&(t\not=-1),\\ j(E_{3,-1,a})=0\end{cases}$
for the $j$-invariants of the elliptic curves from (4.1), (4.2) and (4.3).
This complete the proof of Theorem 1.1. ∎
7. Relation with the reduction type
7.1. Second construction of the pairing
In this subsection, we let $X$ be a smooth projective geometrically connected variety over a field $k$ of dimension $d$. Let ${\operatorname{CH}}^{i}(X)$ denote the Chow group of codimension $i$ cycles on $X$, and put $A_{0}(X):=\ker(\deg:{\operatorname{CH}}^{d}(X)\to\mathbb{Z})$. We shall construct a biadditive pairing
(7.1) $\langle\cdot,\cdot\rangle:{\operatorname{CH}}^{1}(X)_{\operatorname{tors}}\times A_{0}(X)\to k^{\times}\otimes\mathbb{Q}/\mathbb{Z},$
and show that it agrees with (2.1) when $d=1$ (under the identification ${\operatorname{CH}}^{1}(X)=\operatorname{Pic}(X)$). This construction is a key ingredient in the proof of Proposition 7.3 below.
We shall use Bloch’s higher Chow group ${\operatorname{CH}}^{r}(X,i,R)$ with coefficient in a commutative ring $R$. We abbreviate ${\operatorname{CH}}^{r}(X,i):={\operatorname{CH}}^{r}(X,i,\mathbb{Z})$ so that ${\operatorname{CH}}^{r}(X)={\operatorname{CH}}^{r}(X,0)$. We fix $n>0$ and define a biadditive map $\langle\cdot,\cdot\rangle$ as the composition of
(7.2) $\begin{split}\langle\cdot,\cdot\rangle&:{\operatorname{CH}}^{1}(X,1;\mathbb{Z}/n\mathbb{Z})\times{\operatorname{CH}}^{d}(X,0;\mathbb{Z}/n\mathbb{Z})\\ &\to{\operatorname{CH}}^{d+1}(X,1;\mathbb{Z}/n\mathbb{Z})\overset{s_{*}}{\to}{\operatorname{CH}}^{1}(\operatorname{Spec}k,1;\mathbb{Z}/n\mathbb{Z})\cong k^{\times}\otimes\mathbb{Z}/n\mathbb{Z},\end{split}$
where the first map is given by the multiplicative structure of the higher Chow groups, and $s_{*}$ is the push-forward along the structure map $s:X\to\operatorname{Spec}k$. We have an exact sequence and an isomorphism
$0\to{\operatorname{CH}}^{1}(X,1)/n\overset{j}{\to}{\operatorname{CH}}^{1}(X,1;\mathbb{Z}/n\mathbb{Z})\to{\operatorname{CH}}^{1}(X)[n]\to 0,$
${\operatorname{CH}}^{d}(X)/n\cong{\operatorname{CH}}^{d}(X,0;\mathbb{Z}/n\mathbb{Z}).$
We claim that $\langle j(a),b\rangle=0$ for $a\in{\operatorname{CH}}^{1}(X,1)$ and $b\in A_{0}(X):=\ker(\deg:{\operatorname{CH}}^{d}(X)\to\mathbb{Z})$. To show this, we first recall that ${\operatorname{CH}}^{1}(V,1)\cong\mathscr{O}(V)^{\times}$ for any smooth variety $V$. By our assumption on $X$, we obtain isomorphisms
$k^{\times}\cong{\operatorname{CH}}^{1}(\operatorname{Spec}k,1)\overset{s^{*}}{\longrightarrow}{\operatorname{CH}}^{1}(X,1).$
Thus it suffices to show that $\langle j(s^{*}(a)),b\rangle_{X}=a^{\deg(b)}$ for any $a\in{\operatorname{CH}}^{1}(\operatorname{Spec}k,1)$ and $b\in{\operatorname{CH}}^{d}(X)$. We may also suppose $b=[x]$ is the class of a closed point $x\in X$. Write $i_{x}:x=\operatorname{Spec}k(x)\to X$ for the closed immersion. Now the assertion follows from the commutative diagram
${{\operatorname{CH}}^{1}(X,1)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{-\cdot[x]}$$\scriptstyle{i_{x}^{*}}$${{\operatorname{CH}}^{d+1}(X,1)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{s_{*}}$${{\operatorname{CH}}^{1}(\operatorname{Spec}k,1)}$${{\operatorname{CH}}^{1}(\operatorname{Spec}k,1)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(s\circ i_{x})^{*}}$$\scriptstyle{s^{*}}\scriptstyle{\cong}$$\scriptstyle{[k(x):k].}$${{\operatorname{CH}}^{1}(x,1)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i_{x*}}$$\scriptstyle{(s\circ i_{x})_{*}}$
Therefore we obtain an induced biadditive pairing
(7.3) ${\operatorname{CH}}^{1}(X)[n]\times A_{0}(X)/n\to k^{\times}/n,$
and, by taking colimit over $n$, (7.1) as well.
Proposition 7.1.
Suppose $d=1$. Then the two pairings (2.1) and (7.3) are equal to each other, under the canonical identification ${\operatorname{CH}}^{1}(X)\cong\operatorname{Pic}(X)$.
Proof.
Take $a\in\operatorname{Pic}(X)_{\operatorname{tors}}\ (a\not=0)$ and $b\in\operatorname{Pic}^{0}(X)=A_{0}(X)$. We choose $D,E\in\operatorname{Div}(X)$ such that $a=[D],b=[E]$ and $|D|\cap|E|=\emptyset$. We also take $n\in\mathbb{Z}_{>0}$ such that $na=0$ and $f\in k(X)^{\times}$ such that $\operatorname{div}(f)=nD$. Writing $E=\sum_{j}e_{j}Q_{j}$, we have $\langle a,b\rangle=\prod_{j}N_{k(Q_{j})/k}(f(Q_{j}))^{e_{j}}\otimes(1/n)$ under (2.1). We shall prove the same formula for (7.3).
Let $\overline{\Gamma}_{f}\subset X\times\mathbb{P}^{1}$ be the graph of the finite morphism $X\to\mathbb{P}^{1}$ defined by $f$, and $\Gamma_{f}:=\overline{\Gamma}_{f}\cap(X\times\operatorname{\square})$, where $\operatorname{\square}:=\mathbb{P}^{1}\setminus\{1\}$. For $\epsilon\in\{0,\infty\}\in\operatorname{\square}$, denote by $i_{\epsilon}:X\to X\times\operatorname{\square}$ the corresponding closed immersion. Then $\Gamma_{f}$ intersects properly with faces, and its boundary is given by $i_{0}^{*}(\Gamma_{f})-i_{\infty}^{*}(\Gamma_{f})=\operatorname{div}(f)=nD$. Hence $\Gamma_{f}$ defines an element of ${\operatorname{CH}}^{1}(X,1;\mathbb{Z}/n\mathbb{Z})$ whose image in ${\operatorname{CH}}^{1}(X)[n]=\operatorname{Pic}(X)[n]$ agrees with $[D]=a$. The product $\Gamma_{f}]\cdot b\in{\operatorname{CH}}^{2}(X,1;\mathbb{Z}/n\mathbb{Z})$ is represented by the intersection product $\Gamma_{f}\cdot(E\times\operatorname{\square})$ on $X\times\operatorname{\square}$, and its push-forward onto $\operatorname{Spec}k$ by $\sum_{j}e_{j}f(Q_{j})\in\operatorname{Div}(\operatorname{\square})$. By the definition of the isomorphism ${\operatorname{CH}}^{1}(\operatorname{Spec}k,1;\mathbb{Z}/n\mathbb{Z})\cong k^{\times}\otimes\mathbb{Z}/n\mathbb{Z}$ (see [17, p. 183]), this element corresponds to $\prod_{j}N_{k(Q_{j})/k}(f(Q_{j}))^{e_{j}}\otimes(1/n)\in k^{\times}\otimes\mathbb{Z}/n\mathbb{Z}$. This completes the proof. ∎
Definition 7.2.
We define the intrinsic subgroups of $X$ by
${\operatorname{CH}}^{1}(X)_{\operatorname{tors}}^{\operatorname{is}}:=\{a\in{\operatorname{CH}}^{1}(X)_{\operatorname{tors}}\mid\langle a,b\rangle=0\text{ for all }b\in A_{0}(X)_{\operatorname{tors}}\},$
$A_{0}(X)_{\operatorname{tors}}^{\operatorname{is}}:=\{b\in A_{0}(X)_{\operatorname{tors}}\mid\langle a,b\rangle=0\text{ for all }a\in{\operatorname{CH}}^{1}(X)_{\operatorname{tors}}\}.$
7.2. Good reduction
As an application of the second construction of the pairing, we prove that there is some restriction on the possible values of (7.1) if $X$ has good reduction with respect to a discrete valuation.
Proposition 7.3.
Let $v$ be a discrete valuation on $k$, and $O$ its valuation ring. Let $\mathscr{X}\to\operatorname{Spec}O$ be a smooth proper morphism with geometrically connected fibers, and denote by $X$ the generic fiber. Then we have
$\operatorname{Im}(\langle\cdot,\cdot\rangle:{\operatorname{CH}}^{1}(X)_{\operatorname{tors}}\times A_{0}(X)\to k^{\times}\otimes\mathbb{Q}/\mathbb{Z})\subset\operatorname{Im}(O^{\times}\otimes\mathbb{Q}/\mathbb{Z}\to k^{\times}\otimes\mathbb{Q}/\mathbb{Z}).$
Proof.
We fix $n>0$ and prove
$\operatorname{Im}(\langle\cdot,\cdot\rangle:{\operatorname{CH}}^{1}(X)[n]\times A_{0}(X)/n\to k^{\times}/n)\subset\operatorname{Im}(O^{\times}/n\to k^{\times}/n).$
We use Bloch’s higher Chow group for schemes over a discrete valuation ring, for which we refer to [8]. Take $a\in{\operatorname{CH}}^{1}(X)[n]$ and $b\in A_{0}(X)/n$. We shall show $\langle a,b\rangle\in O^{\times}/n$. Let $d:=\dim X$. Since ${\operatorname{CH}}^{1}(\mathscr{X})\overset{\cong}{\to}{\operatorname{CH}}^{1}(X)$ is an isomorphism, the map $\phi$ in the diagram
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${O^{\times}/n\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\operatorname{CH}}^{1}(\mathscr{X},1;\mathbb{Z}/n\mathbb{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$${{\operatorname{CH}}^{1}(\mathscr{X})[n]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${k^{\times}/n\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\operatorname{CH}}^{1}(X,1;\mathbb{Z}/n\mathbb{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\operatorname{CH}}^{1}(X)[n]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$
[…]
$\psi:{\operatorname{CH}}^{d}(\mathscr{X},1,\mathbb{Z}/n\mathbb{Z})\twoheadrightarrow{\operatorname{CH}}^{d}(X,1,\mathbb{Z}/n\mathbb{Z})\twoheadrightarrow{\operatorname{CH}}^{d}(X)/n.$
[…]
${{\operatorname{CH}}^{1}(\mathscr{X},1;\mathbb{Z}/n\mathbb{Z})\times{\operatorname{CH}}^{d}(\mathscr{X})/n\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma}$${{\operatorname{CH}}^{1}(X,1;\mathbb{Z}/n\mathbb{Z})\times{\operatorname{CH}}^{d}(X)/n\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\operatorname{CH}}^{d+1}(\mathscr{X},1;\mathbb{Z}/n\mathbb{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\operatorname{CH}}^{d+1}(X,1;\mathbb{Z}/n\mathbb{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\operatorname{CH}}^{1}(\operatorname{Spec}O,1;\mathbb{Z}/n\mathbb{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${{\operatorname{CH}}^{1}(\operatorname{Spec}k,1;\mathbb{Z}/n\mathbb{Z})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${O^{\times}/n\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$${k^{\times}/n,}$
we find that $\langle a,b\rangle=i(\gamma(\widetilde{a},\widetilde{b}))\in O^{\times}/n$. We are done. ∎
Example 7.4.
Let $p$ be a prime and $k=\mathbb{Q}_{p}$. Then ${\operatorname{CH}}^{1}(X)_{\operatorname{tors}}^{\operatorname{is}}$ (resp. $A_{0}(X)_{\operatorname{tors}}^{\operatorname{is}}$) contains all elements of ${\operatorname{CH}}^{1}(X)_{\operatorname{tors}}$ (resp. $A_{0}(X)_{\operatorname{tors}}$) having order prime to $p$. This follows from Proposition 7.3, since $\mathbb{Z}_{p}^{\times}\otimes\mathbb{Q}/\mathbb{Z}\cong\mathbb{Q}_{p}/\mathbb{Z}_{p}$ is $p$-primary torsion. In particular, we have $E(\mathbb{Q}_{p})^{\operatorname{is}}_{\operatorname{tors}}=E(\mathbb{Q}_{p})_{\operatorname{tors}}$ if $p\geq 5$ and if $E$ is an elliptic curve with good supersingular reduction (so that $|E(\mathbb{Q}_{p})_{\operatorname{tors}}|$ is a divisor of $p+1$).
The following corollary is an immediate consequence of Proposition 7.3.
Corollary 7.5.
Suppose that $k$ is a number field and that $X$ has good reduction outside a finite set $S$ of places of $k$ containing all archimedean places. Let $O_{k,S}$ be the ring of $S$-integers. Then we have for any $D,E\in\operatorname{Div}_{t}(X)$
$\langle D,E\rangle\in\operatorname{Im}(O_{k,S}^{\times}\otimes\mathbb{Q}/\mathbb{Z}\to k^{\times}\otimes\mathbb{Q}/\mathbb{Z}).$
Example 7.6.
Let $E$ be an elliptic curve over $\mathbb{Q}$ and let $p_{1},\dots,p_{n}$ be all the primes at which $E$ has bad reduction. Then we have for any $P,Q\in E(\mathbb{Q})_{\operatorname{tors}}$
$\langle P,Q\rangle\in\left(\{\pm 1\}\cdot\bigoplus_{i=1}^{n}p_{i}^{\mathbb{Z}}\right)\otimes\mathbb{Q}/\mathbb{Z}\subset\mathbb{Q}^{\times}\otimes\mathbb{Q}/\mathbb{Z}.$
7.3. Tate curves
In this subsection, we study the intrinsic subgroup of an elliptic curve with split multiplicative reduction.
Assume that $k$ is complete with respect to a normalized discrete valuation $v:k^{\times}\to\mathbb{Z}$. We take $q\in k^{\times}$ such that $n:=v(q)>0$, and let $E_{q}:=\mathbb{G}_{m}/q^{\mathbb{Z}}$ be the Tate curve with parameter $q$. We write $a]\in E_{q}(k)$ for the point corresponding to the class of $a\in k^{\times}$ under the canonical isomorphism $E_{q}(k)\cong k^{\times}/q^{\mathbb{Z}}$. We have exact sequences
(7.4) $0\to O_{k}^{\times}\overset{[\cdot]}{\to}E_{q}(k)\overset{v}{\to}\mathbb{Z}/n\mathbb{Z}\to 0,\quad\text{and}\quad 0\to\mu(k)\overset{[\cdot]}{\to}E_{q}(k)_{\operatorname{tors}}\overset{v}{\to}\mathbb{Z}/n\mathbb{Z},$
where $O_{k}$ is the valuation ring of $k$. We take $m>0$ and $a\in k^{\times}$ such that $a]\in E_{q}(k)[m]$. We then have $a^{m}=q^{\widetilde{s}_{m}(a)}$ for some $\widetilde{s}_{m}(a)\in\mathbb{Z}$, which is necessarily given by $\widetilde{s}_{m}(a)=v(a)m/n$. Thus we obtain a well-defined map
$s_{m}:E_{q}(k)[m]\to\mathbb{Z}/m\mathbb{Z},\qquad s_{m}([a])=\widetilde{s}_{m}(a)\bmod m\in\mathbb{Z}/m\mathbb{Z}.$
This fits in an exact sequence
(7.5) $0\to\mu_{m}(k)\overset{[\cdot]}{\to}E_{q}(k)[m]\overset{s_{m}}{\to}\mathbb{Z}/m\mathbb{Z}\overset{\delta_{m}}{\to}k^{\times}/(k^{\times})^{m},$
where $\delta_{m}$ is given by $\delta_{m}(1)=q\cdot(k^{\times})^{m}$.
Proposition 7.7.
  1. (1)
    Let $a,b\in k^{\times}$ and $m>0$. If $a],[b]\in E_{q}(k)[m]$, we have
    $\langle[a],[b]\rangle=a^{-s_{m}([b])}\otimes\frac{1}{m}=b^{-s_{m}([a])}\otimes\frac{1}{m}=q^{-s_{m}([a])s_{m}([b])}\otimes\frac{1}{m^{2}}.$
  2. (2)
    We have $E_{q}(k)_{\operatorname{tors}}^{\operatorname{is}}=\{[\zeta]\mid\zeta\in\mu(k)\}$.
Proof.
(1) Put $s:=\widetilde{s}_{m}(a),s^{\prime}:=\widetilde{s}_{m}(b)\in\mathbb{Z}$. The last two equalities easily follow from $a^{m}=q^{s},\ b^{m}=q^{s^{\prime}}$. We show the first equality. It is obvious if $[a]=[1]$, hence we assume $a]\not=[1]$ in what follows. We use the $p$-adic theta function
$\theta(u):=(1-u)\prod_{i\geq 1}\frac{(1-q^{i}u)(1-q^{i}u^{-1})}{(1-q^{i})}^{2}$
which enjoys the fundamental equality (see [15, V, Proposition 3.2]):
(7.6) $u\theta(qu)=\theta(u).$
It follows from this formula (and $a^{m}=q^{s}$) that
$f(u):=\frac{\theta(u/a)^{m}}{\theta(u)^{m-1}\theta(u/q^{s})}$
satisfies $f(qu)=f(u)$. Hence we have $f(u)\in k(E_{q})^{\times}$ and $\operatorname{div}(f)=m([a]-[1])$. We claim
(7.7) ${\operatorname{lc}}_{[b]}(f,m)=q^{s(s-1)/2}b^{-s}\otimes\frac{1}{m}$
for any $b\in k^{\times}$. Indeed, when $[b]=[1]$ this is seen as
${\operatorname{lc}}_{[1]}(f,m)$ $=\left(\frac{f(u)}{\theta(u)^{m}}\right)|_{u=1}\otimes\frac{1}{m}=\theta(1/a)^{m}\cdot\left(\frac{\theta(u)}{\theta(u/q^{s})}\right)|_{u=1}\otimes\frac{1}{m}$
$=\theta(1/a)^{m}\cdot(-1)^{s}q^{s(s-1)/2}\otimes\frac{1}{m}=q^{s(s-1)/2}\otimes\frac{1}{m},$
where we used (7.6) again. When $[b]=[a]$ we proceed similarly:
${\operatorname{lc}}_{[a]}(f,m)$ $=\left(\frac{f(u)}{\theta(u/a)^{m}}\right)|_{u=a}\otimes\frac{1}{m}=\frac{1}{\theta(a)^{m-1}\theta(a/q^{s})}\otimes\frac{1}{m}$
$=\theta(a)^{-m}\cdot(-1)^{s}q^{s(s-1)/2}a^{-s}\otimes\frac{1}{m}=q^{s(s-1)/2}a^{-s}\otimes\frac{1}{m}.$
Finally, it is easier when $b]\not=[1],[a]$:
${\operatorname{lc}}_{[b]}(f,m)$ $=f(b)\otimes\frac{1}{m}=\left(\frac{\theta(b/a)}{\theta(b)}\right)^{m}\frac{\theta(b)}{\theta(b/q^{s})}\otimes\frac{1}{m}=q^{s(s-1)/2}b^{-s}\otimes\frac{1}{m}.$
We have shown (7.7), from which (1) readily follows.
(2) Since $s_{m}([\zeta])=0$ for any $m>0$ and $\zeta\in\mu_{m}(k)$, the right hand side $\mu(k)]$ of (2) is contained in $E_{q}(k)_{\operatorname{tors}}^{\operatorname{is}}$ by (1). Hence $\langle\cdot,\cdot\rangle$ factors through a symmetric biadditive pairing on $T:=E_{q}(k)_{\operatorname{tors}}^{\operatorname{is}}/[\mu(k)]$. From (7.4) we find that $T$ is a finite cyclic group and $m_{T}:=|T|$ is divisible by $n$. Take $a\in k^{\times}$ such that the class of $a]\in E_{q}(k)_{\operatorname{tors}}$ generates $T$, and let $m$ be the order of $[a]$ in $E_{q}(k)_{\operatorname{tors}}$ (which can be strictly larger than $m_{T}$). Note that we have $m/m_{T}=\min\{\nu>0\mid q^{\nu}\in(k^{\times})^{m}\}$ and $a^{m}=\zeta q^{cm/m_{T}}$ for some $\zeta\in\mu_{m}(k)$ and $c\in\mathbb{Z}$ with $c$ coprime to $m_{T}$. Now (1) shows that $\langle[a],[a]\rangle=a\otimes(-c^{2}/m_{T})\in k^{\times}\otimes\mathbb{Q}/\mathbb{Z}$, which has order precisely $m_{T}$ as seen from the exact sequence
$\mu(k)\overset{m}{\to}\mu(k)\to k^{\times}/(k^{\times})^{m}\to k^{\times}\otimes\mathbb{Q}/\mathbb{Z}.$
Thus $a]^{\nu}$ belongs to $E_{q}(k)_{\operatorname{tors}}^{\operatorname{is}}$ if and only if $\nu$ is a multiple of $m_{T}$. We are done. ∎
Corollary 7.8.
Let $E$ be an elliptic curve over a number field $k$, and $P\in E(k)_{\operatorname{tors}}^{\operatorname{is}}$ an element of the intrinsic subgroup of order $m>0$. Let $v$ be a finite place of $k$ such that the completion $k_{v}$ of $k$ at $v$ satisfies $\mu_{m}(k_{v})=\{1\}$. If $E$ has split multiplicative reduction of Kodaira type $I_{n}$ at $v$, then $n$ is divisible by $m^{2}$.
Proof.
The assumption implies that the base change of $E$ to $k_{v}$ is isomorphic to the Tate curve $E_{q}=\mathbb{G}_{m}/q^{\mathbb{Z}}$ for some $q\in k_{v}$ such that $v(q)=n$ ([15, Chapter V, Theorem 5.3]) Let us write the image of $P$ in $E_{q}(k_{v})$ as $[a]$ with $a\in k_{v}$. Since $\mu_{m}(k_{v})=\{1\}$, we find from (7.5) that $E_{q}(k_{v})[m]$ is a cyclic group of order $m$ generated by $[a]$. In particular, $s:=s_{m}([a])$ is invertible in $\mathbb{Z}/m\mathbb{Z}$. We have $\langle[a],[a]\rangle=0$ since $P\in E(k)_{\operatorname{tors}}^{\operatorname{is}}$. On the other hand, Proposition 7.7 (1) shows $\langle[a],[a]\rangle=q^{-s^{2}}\otimes(1/m^{2})$. By looking at its image under the map $v\otimes{\operatorname{id}}_{\mathbb{Q}/\mathbb{Z}}:k_{v}^{\times}\otimes\mathbb{Q}/\mathbb{Z}\to\mathbb{Q}/\mathbb{Z}$, we conclude that $s^{2}n/m^{2}=0$ holds in $\mathbb{Q}/\mathbb{Z}$, that is, $m^{2}$ divides $n$. ∎
Example 7.9.
Let $E$ an elliptic curve over $\mathbb{Q}$ such that $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}$ contains an element of order $m>0$. Let $p$ be an odd prime such that p1p-1 is prime to $m$. If $E$ has split multiplicative reduction of Kodaira type $I_{n}$ at $p$, then $n$ is divisible by $m^{2}$.

### v2

@@
 Abstract.
-We define and study a biadditive symmetric (not necessarily perfect) pairing on the torsion part $\mathrm{Pic}(X)_{{\operatorname{tors}}}$ of the Picard group of a smooth projective curve $X$ over a field $k$ with values in $k^{\times}\otimes\mathbb{Q}/\mathbb{Z}$.
+We define and study a biadditive symmetric (not necessarily perfect) pairing on the torsion part $\mathrm{Pic}(X)_{\mathrm{tors}}$ of the Picard group of a smooth projective curve $X$ over a field $k$ with values in $k^{\times}\otimes\mathbb{Q}/\mathbb{Z}$.
 We call its kernel the intrinsic subgroup of $X$.
 It turns out that some information on the reduction type of $X$ can be read off from the intrinsic subgroup.
-Mazur’s torsion theorem says that there are exactly 15 isomorphism classes of abelian groups that appear as the rational torsion points of an elliptic curve $X$ over $\mathbb{Q}$ (identified with $\mathrm{Pic}(X)_{\operatorname{tors}}$).
+Mazur’s torsion theorem says that there are exactly 15 isomorphism classes of abelian groups that appear as the rational torsion points of an elliptic curve $X$ over $\mathbb{Q}$ (identified with $\mathrm{Pic}(X)_{\mathrm{Tor}}$).
 We refine this result by determining which subgroups of those 15 groups appear as the intrinsic subgroups.
@@
 Picard group of a curve, torsion rational points of an elliptic curve.
-2010 Mathematics Subject Classification:
-11G18, 14G05, 14G35
 2010 Mathematics Subject Classification:
@@
 Next, if $\epsilon=+$ or $M$ is odd, we can apply Remark 5.2 (1) and Corollary 5.7 to deduce $Y_{1}(N)_{M}^{\epsilon}(\mathbb{Q})=\emptyset$.
-This covers the cases $(M,N)\in\{(7,7),(10,5),(12,3)\}$ with $\epsilon\in\{+,-\}$, and $(M,N)\in\{(6,6),(8,4),(10,2),(12,2)\}$ with $\epsilon=$-.
+This covers the cases $(N,M)\in\{(7,7),(10,5),(12,3)\}$ with $\epsilon\in\{+,-\}$, and $(N,M)\in\{(6,6),(8,4),(10,2),(12,2)\}$ with $\epsilon=$-.
 This leaves us with the cases $(N,M)\in\{(6,6),(8,4),(9,3),(10,2),(12,2)\}$ with $\epsilon=$-.
@@
 Case of $A=\mathbb{Z}/4\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$
-Let $(E,P,Q)\in Y_{1}^{0}(4,2)(\mathbb{Q})$, and suppose first $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}$ is non-trivial.
-Then we must have $E(\mathbb{Q})_{\operatorname{tors}}\cong A$ (because we have shown $E(\mathbb{Q})_{\operatorname{tors}}\cong\mathbb{Z}/8\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$ implies $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=0$), and then $|E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}|=2$ (because we have shown $|E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}|\not=4,8$).
-Thus we have
-  • $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=\langle 2P\rangle\Leftrightarrow(E,P,Q)\in\operatorname{Im}(Y_{1}^{0}(4,2)_{(2,1,1)}^{\pm}(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q}))$,
-  • $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=\langle Q\rangle\Leftrightarrow(E,P,Q)\in\operatorname{Im}(Y_{1}^{0}(4,2)_{(1,2,2)}^{\pm}(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q}))$.
-Since the components of $Y_{1}^{0}(N)_{(2,1,1)}^{\pm}$ and $Y_{1}^{0}(N)_{(1,2,2)}^{\pm}$ are isomorphic to $\mathbb{P}^{1}$ by Proposition 5.15, we conclude by Proposition 6.2 that there are infinitely many such $(E,P,Q)$ in both cases.
-Finally, we have $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=0$ precisely when neither of the above conditions are met and $(E,P,Q)\not\in\operatorname{Im}(Y_{1}^{0}(8,2)(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q}))$.
-Again we conclude by Proposition 6.2 that there are infinitely many such $(E,P,Q)$.
+Let us introduce three subsets of $Y_{1}^{0}(4,2)(\mathbb{Q})$:
+$J_{1}:=$ $\operatorname{Im}(Y_{1}^{0}(8,2)(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q})),$
+$J_{2}:=$ $\operatorname{Im}(Y_{1}^{0}(4,2)^{\pm}_{(2,1,1)}(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q})),$
+$J_{3}:=$ $\operatorname{Im}(Y_{1}^{0}(4,2)^{\pm}_{(1,2,2)}(\mathbb{Q})\to Y_{1}^{0}(4,2)(\mathbb{Q})).$
+Note that $J_{1}\cap J_{3}=\emptyset$.
+Indeed, since $Y_{1}^{0}(8,2)_{(1,2,1)}^{\pm}(\mathbb{Q})=\emptyset$ by Proposition 5.15, we have $\langle Q,Q\rangle\not=0$ for any $(E,P^{\prime},Q)\in Y_{1}^{0}(8,2)^{\pm}(\mathbb{Q})$, which implies the claim.
+For $(E,P,Q)\in Y_{1}^{0}(4,2)(\mathbb{Q})$, we have $E(\mathbb{Q})_{\operatorname{tors}}=\langle P,Q\rangle\Leftrightarrow(E,P,Q)\not\in J_{1}$.
+Suppose this holds.
+Then we know that $|E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}|\leq 2$ (by the non-existence part).
+Therefore we have the equivalences
+  • $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=\langle 2P\rangle\Leftrightarrow(E,P,Q)\in J_{2}$,
+  • $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=\langle Q\rangle\Leftrightarrow(E,P,Q)\in J_{3}$,
+  • $E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}=0\Leftrightarrow(E,P,Q)\not\in J_{2}\cup J_{3}$.
+We conclude the existence part of Theorem 1.1 in this case by Proposition 6.2.
 Case of $A=\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$
 We need a special treatment for this case. Consider the following subsets of $Y_{1}(1)(\mathbb{Q})$.
-$J:=$ $\operatorname{Im}(Y_{1}(2,2)(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
-$J_{1}:=$ $\operatorname{Im}(Y_{1}(4,2)(\mathbb{Q})\sqcup Y_{1}(6,2)(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
-$J_{2}:=$ $\operatorname{Im}(Y_{1}(4,2)_{(2,1,1)}^{\pm}(\mathbb{Q})\sqcup Y_{1}(4,2)_{(1,2,2)}^{\pm}(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
-$J_{3}:=$ $\{j(E)\in Y_{1}(1)(\mathbb{Q})\mid(E,P,Q)\in Y_{1}(2,2)(\mathbb{Q}),\ E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}\not=0\}.$
+$J:=$ $\operatorname{Im}(Y_{1}^{0}(2,2)(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
+$J_{1}:=$ $\operatorname{Im}(Y_{1}^{0}(4,2)(\mathbb{Q})\sqcup Y_{1}^{0}(6,2)(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
+$J_{2}:=$ $\operatorname{Im}(Y_{1}^{0}(4,2)_{(2,1,1)}^{\pm}(\mathbb{Q})\sqcup Y_{1}^{0}(4,2)_{(1,2,2)}^{\pm}(\mathbb{Q})\to Y_{1}(1)(\mathbb{Q})),$
+$J_{3}:=$ $\{j(E)\in Y_{1}(1)(\mathbb{Q})\mid(E,P,Q)\in Y_{1}^{0}(2,2)(\mathbb{Q}),\ E(\mathbb{Q})_{\operatorname{tors}}^{\operatorname{is}}\not=0\}.$
 By Corollary 5.14 and the (proved) non-existence result, we obtain